A Complete Bibliography of Publications in Statistics and Probability Letters: 1980–1989

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
      WWW URL: http://www.math.utah.edu/~beebe/

14 October 2017 Version 1.03

Title word cross-reference

(θ_{1} + θ_{2} + θ_{3})/θ_{4} [BB89]. 1 < r < 2 [CS87a]. 2 × 2 [Bök88, Wal85, WD88]. β [CM84]. β = 2 [GG89]. C(α) [Akr88]. C_{α} [BW89]. χ^{2} [Ray88, CF83, Sve85]. d [Shi88]. F [CF83, GS87]. k [Alb88, Flu86a, Lin88, Lin89, Par88, Sca85]. κ [HA87, PAT88]. L [Koe84]. L^{2} [Lar84, PA88]. l^{∞} [Mis89]. L_{1} [Akr82, FF89]. L_{2} [Dev89a, Lec85, Raf87]. M [AJ89, Cha86, JKV88, FM87]. R^{k} [Mee86]. N [Pap84, Kem88, Par88, Sca85]. N = 11(1)18 [Fra88]. P [Cha86]. P (Y < X) [YM84]. Π [AA88]. R [RS88, CM88b]. sup |S_{n}|/n^{1}/r [CS87a]. t [CF83]. T_{0}^{2} [Ksh89]. U [Deh89, Kok87, Ses88]. W [SH87]. X_{n} = c min(X_{n−1}, Y_{n}) [AH89a]. Z_{p} [Cha89].

-content [CM84]. -cross-validated [Dev89a]. -dimensional [Sca85]. -estimates [Koe84]. -estimators [JKV88]. -extensible [Sca85]. -group [FM87]. -means [CM88b]. -norm [FF89, Akr82]. -optimal [Lec85]. -quantiles [AJ89]. -sample [Alb88]. -statistic [Ses88]. -statistics [Deh89, Kok87]. -stochastic [PA88]. -test [Sve85]. -tests [BW89]. -variation [AA88].

1988/1989 [Ano89a].

7 [Ano89d].

Abel [Hor85]. absolute [HW88, MS87b]. absolutely [Ebe84]. accelerated [GM85, JC89]. accuracy [HM89]. ace [Ket89]. Acknowledgement [Ros84]. adaptation [LY89]. adaptive [Fab88]. additional [Kan84]. Adjacent [DG88]. adjustment [Sch89a]. Admissibility [CS87b, Ste86]. Admissible [VM83, Yao89]. against [CD88, KD85]. Aggregation [FRT84]. aging [KD85]. agreement [Agr89]. aids [GJ88a]. Akritas [Ano86d]. Algebraic [Dal87]. algebras [MW87]. algorithm [CF83, DG88, Fu89, HS89, Hat86, Ket89, OI83, ZR86]. aligned [KS85, Sen84a]. allocation [FFT84, Har89]. Almost [Mas87, MS88, ER87, How83]. alternatives [CD88]. among [AH89a]. analogue [Sim83]. analysis [AC82, CKS88, DM84, Fin87, FM87, GM85, IO85, Iha87, JC89, KPW84, LW82, LC82, MS87b, Sha89a, Tsu88a, Wan85]. ancillarity [Bha87]. ANOVA [WK83]. antibodies [GJ88a]. application [AA89, GJ88a, KPW84, LM87, Wan85, Wel84, Wit85, YR88]. Applications [Bal89, Geo84, Pet89, Sto83, dP84, dP87]. applied [JH84]. approach [AB86, AT85, Gou89b, MS85]. approximate [JH84, Rei86]. Approximation [Hor85, Ahm85, Fra88, GS87, Hor84, Kor87a, Smo85, Ste88]. Approximations [CH86, CH89c, Hou85]. arbitrarily [GC86]. arbitrary [MS88]. argument [Bha89]. Arima [LR89]. arithmetic [Sen87]. ARMA [AP86]. arrays [Deh89, Edl84]. assessing [Sri84, SH87]. assets [AC82]. associated [Dab85, PX89]. association [AA89, Bec89, Bök88, Tsu88b]. asymmetric [AH89b]. Asymptotic [BBW89, Bra83, CR89, Geo88, Hus88, KL85b, Krz83, Mai88, Mar83, Mie87a, MS84, Rao85, Sar86, Akr88, BLE87, BM87, Ber89, DM84, Fab88, Flu86b, Gou89a, HK88, Kha86, KV85, Kni89, Rou89b, Sha89a, SB88, Yin89]. Asymptotically [Wel88, KL85a]. asymptotics [Ath87]. Attainable [Roo84, MM88b]. attaining [Raf87]. attraction [Mee86, Tak87]. attribute [AK83]. Author [Ano83, Ano84a, Ano85a, Ano87a, Ano88a, Ano89a, Ano89b]. autocorrelations [LR89]. autocovariances [LR89]. autoregressive [CD82, Har84a, Le 88, Par82]. auxiliary [Gup83, Hun85]. average [Gri88, LY89].

Backward [Tay82]. Bahadur [Baj87, FK84]. balance [And86, CK87, Kat89]. balanced [Afs88a, Joh85, Kag88, SJ88, Sin89, Tal86, WK83]. Banach [Hol89, Rhe84]. bands [Mie87a, Pie87]. Bandwidth [Chi89]. based [Abd89, AH89b, BLE87, CR89, CG84, HS86, HR89a, KL85a, Mie87a, NH89, Pes82, Pes84b, Rao84b, SH87, TH87, Wel84]. Bayes [And84, Kuo86, Li85, Mar89, Wu89]. Bayesian [AB86, CKS88, FM87, JC89, MS87a, MS85, Sin86]. be [Bal89, CHR86, Dan88, How83]. Bechhofer [Lev84]. behavior [AAO89, DLS86, Edl84, ER87, Gou89a, Ros83, Ros84]. behaviour [Ber89, LR89, Mai88]. Bellman [Tay82]. belongs [Tak87]. Berkes [Dab82]. Bernoulli [Sto83]. Bessel [Ath87]. best [CH89c]. beta [TG84]. better [BE86]. Between [HR89b, CK87, CR86, CH89b, CHR86, Dav86, Eub82, GM85, HM87b, Kap88, RS89, Saw82, TC88]. Bhattacharya [CP86]. Bias [Tsa89, CG84, SB88]. biased [Kat89, Sen87]. BIB [CK87]. binary [GK89, How83, Kag88, Pes82, Pes84b]. binary-type [Pes82, Pes84b]. binomial [GS87, Koc89, Lin88, Lin89]. Birnbaum [YM84]. birth [Tan86]. birth-death [Tan86]. bits [The87, TC88]. Bivariate [Arn87, Bec89, BR88, DdO89, GJ88b, KV85, Koc89, Kor87b, O’N85, dO89]. Block [Yan88, PT86, SJ88, Sin89]. BLUE [Bal89]. Board [Ano82a, Ano85b, Ano87b, Ano88b]. Bonferroni [Hop85, TX89]. Bootstrap [BB88, Ray89, Ath83, BG88b, HM89, Par85]. bootstrapping [LB83]. bootstraps [TRF84]. both [CM84]. bound [Dev89b, Feu88, Fu83, FK84, JH84, Rot83, YM84]. Boundary [RC87, Rus88]. bounded [SYM86, Wu89]. bounding [Tho89, Vit88]. Bounds [CP86, CR86, Hoo89, YR88, AB88, CP85, CP89, Dal87, Hop85, MM88b, Pap88, Pes82, Sca85, SYM86]. box [Hol89]. branching [Tay82]. breakdown [Cha86]. Brownian [CHR86, KS87a, Sim83]. busy [SH89].

Calculating [JP84]. calculus [Sha89a]. calibration [Sun85]. can [How83]. cannot [Kor89]. Canonical [MN88]. carcinogenesis [Ser84, SKM86]. Carlo [Fin87]. carriers [BGM^{+}86]. case [Geo88, WJ88]. cases [BP85]. categorical [Bec89]. categories [Sim87a]. category [Tsu88b]. Cauchy [Pap88, Phi89]. caution [Bra89]. cell [Lec85]. cells [Ray88]. censored [GC86, Kao85, Kor87b, PT88, Tsa85]. censoring [Ber83, LP87, Liu86, Mie87a, Sta86]. censorship [Alb88]. central [Zel88, Joh88, Pfl83, Rhe84, Rhe86, iYT87]. certain [BH84, Hor84, SYM86]. certainly [How83]. certificates [GL89]. chain [AT89, AC82, SS83]. chains [Ath85, Bur89, Joh88]. change [Yao88]. change-points [Yao88]. characteristic [Dev84, Kam82, Wel84, Wel86]. Characterization [Dan89, Gzy87, KM87a, KS87b, NH89, Abd85, AH89a, Bal87, FF89, FS85, Gat88, GR89, HC89, JH84, Kor85, KS88, RS86a, Wes87, Yao89]. Characterizations [Cac83, Cac87, CP89, EN84, Lus89, TL89, BT84, YM89]. chemical [CK87]. Chernoff [Akr84, Ano86d, KS87b]. chi [BH84, BG88a, BH86a, Bon89, Sch87]. chi-square [BG88a, BH86a, Bon89]. chi-squared [Sch87]. chi-squaredness [BH84]. choice [Gol88, MS87a]. choosing [Kap88]. circle [LB83]. Circular [Afs88a, Ber89]. circumference [LB83]. class [CD88, Cha85a, Kha86, KM87a, KPW84, Lin89, Raf87, Ruk83, SS84, Sve85, Tsa85, Wan85]. classes [AEN86, EN84, O’B82, Ste86]. classification [Kag85, Per82]. Closed [KL85a]. closure [EN84]. cluster [GG89, Hun87]. clustered [RP85]. clustering [Eve88]. clusters [Gup82]. Clusure [AEN86]. coefficient [FM87, NY88]. Coefficients [MS87b, SS86]. collapsibility [Dav87]. collapsible [Dav86]. column [SGS87]. combine [Wal85]. comment [Wan85]. common [Kan84]. comparing [Ber83]. Comparison [HR89a, Flu86b, IO85, SM87, TRF84]. comparisons [Pes84b]. competing [YR88]. Complete [Deh89, HR89b]. complex [Fin87, Le 88]. complex-valued [Le 88]. component [Khu87, MS88, Sch87, Sch89a]. components [Bue88, CS88, MG85, Par88, TBGG89]. compound [Ste88]. compounding [Sze88]. computational [DM84]. computing [Fu89]. concentrated [Riv86]. concept [BSS85]. concerning [Eva89, Mie87b, Sch89a, Wef85]. concordance [PS83]. condition [Bal89]. Conditional [MS86, BS84, BBW89, Mie87b, Sam89]. conditionally [CS89, HK88]. conditionals [Arn87]. conditioning [CT82]. Conditions [MS87a, Bra89, Dan88, Rhe86, Rou89a, Rou89b]. Confidence [BB89, HS86, BB88, HM89, Jey85, Mie87a, Pie87, Ray89, SM87, TBGG89]. conforming [Kan84]. conjugate [AA88, Ath86, MFP89]. connected [Sta89]. Connectedness [DPV84]. connections [Par85]. Consistency [KV85, Abd89, CM88b, Dev89a, GP85, Kou87, MMR88, Mie87b]. Consistent [Rou89a]. constants [Tak87]. constraints [Bon89]. constructed [Arn88]. Constructing [Joh85, Ray88]. Construction [Kan84]. contagious [Wil88]. contaminated [CKS88]. content [CM84]. contiguity [Akr82]. contingency [BH86a, Dav86, Dav87, Lee88, Sim87a]. Continuity [MW87, GK89, Pla88]. continuous [tHK84, OR89]. contoured [KM87a]. control [SKM86]. controlled [Sun85]. controlling [CM84]. Convergence [Ric86, RR83, Bur89, Bur85, CGN83, Deh89, Ebe84, Fra88, FK84, Gou89b, IR87, Kok87, LP87, Mas87, MM87, Raf87, Rao84a, Rhe84, Roo84, RC87, Ver88]. convex [And84, Kor89]. convexity [BP87]. Convolution [Ano89d, MM88a, EN84, Pla88]. convolutions [GK89]. Cook [dR88]. coordinate [RE89]. coordinate-free [RE89]. coordinates [LB83]. correlated [Chi89, MS86, MS87c]. correlation [KPW84, NY88, PS83, Sam84, Wan85]. cost [Pes82, PT86]. counterexample [Wef85]. counterpart [Kor87a]. counterstrategy [Bro82]. covariance [Bon86, CT82, Flu86b, Flu86a, Hol87, Kni89, Kub89]. covariances [Cha85b, JJ87]. covariates [Mat84]. covariogram [CG84]. coverage [Ald89]. Cox [Yan89]. Cramér [JKV88, MM88b]. Cramér-type [JKV88]. credibility [Bro82]. criteria [Abo88, Pet86]. criterion [God87b, Saw82, Yao88]. critical [CF83]. cross [Dev89a, HS89, Lee88]. cross-product [Lee88]. cross-validation [HS89]. crossing [RC87]. crude [YR88]. curve [Par83a]. curved [Hou85]. cusum [Hus88]. cycle [Li87]. cyclic [Li87]. cylindrical [OT87].

data [BS84, Bec89, CKS88, Eve88, GC86, GJ88b, Gup82, HR89a, Kor87b, PT88, SR84, WK83]. death [Tan86]. decomposability [Sze88]. decomposition [KS87a, WK83]. define [Ebr85]. defined [Bar87]. definite [Mis89, Ric85]. definition [CJT87]. degenerate [Kok87]. degree [GG89]. delete [Shi88]. delete- [Shi88]. deletion [MG88]. densities [Len87, Mie87a, Par82]. Density [CM88a, Rao84b, Abd89, Bab86, Bra83, CR89, Dev89a, Fal86, GP85, HM87a, HW88, Har84b, IR87, KV85, MN88, Mas87, Ryc87, Sch89b, Sin87, TS89]. Dependence [CKZ86, BSS85, HR89b, Rou89a, Rou89b, dO89]. dependent [Bir88, Dab82, DT88, Hoo89, Mis89, Sar86, Ser84, iYT87]. deposit [GL89]. derivation [Akr88, Sim83]. derivative [Wig85]. derivatives [Geo84, HM87a, Pet89, Sin87]. Descriptive [Oja83]. design [FB86, Geo88, MMR88, Rou89a]. designs [Afs86, Afs87, Afs88a, Afs88b, CZ84, CK87, DPV84, HM87b, Joh85, Kag85, Kag88, Kat89, MM84, MG88, Mül84, PT86, SGS87, SJ88, Sin89]. destructive [AK83]. determinant [And89, WD88]. Determination [CM84, MS87b]. determined [Str89]. Determining [O’B82, Fin87, Gup82, Sim87b]. develop [LC82]. deviation [Fu83, JKV88, MS87b]. deviations [HH88, Led87, RR83]. diagnostic [Kao85, dR88]. Diagnostics [EG86]. diagonal [Yan88]. diagonals [Agr83]. diagonals-parameter [Agr83]. difference [CR86, CHR86, HM87b, Lee88]. different [FFT84, GJ88a]. differentiable [Par83b]. differential [Dab89, Nua87, SS86, Vit88]. diffusions [Kor89]. digits [How83]. dimensional [AB89, Ald89, HH88, Lus89, Ric85, Sca85]. direction [Kor87a]. directional [JP88, Wig85]. Dirichlet [Kuo86]. disclosure [Alb84]. discrete [Ber83, BT84, Cac83, Cac87, CS87b, Dan88, Dan89, KS87b, NH89, Sim83, Sve85, TRF84]. discriminant [CR89, Krz83]. discriminating [tHK84]. Dispersive [Bar86, AK88, Bar87]. distance [Cla89, HW88, Kou85, Lar84]. distances [BH86a]. distributed [BB86, Hor85, WD88]. Distribution [BB87, AB86, AJZ89, Ath86, Bal87, Bec89, BG88a, Bru88, CD88, Cha85a, CM84, Dal88, Fam87, Gaj87, Gat88, GS87, Har84a, JJ84, Kam82, Kea84, Koc89, Krz83, Li87, MS87a, MB87, Mut88, NH89, OR89, O’N85, PX86, PX89, Pap84, PGP83, RS86a, RS86b, Ric83, Riv86, Sam83, Sca85, Ses88, Sin86, Sta86, Sta87a, SH89, Sti82, Tak87, TG84, The82]. distribution-free [CD88]. distributions [Abd85, AAO89, Arn87, AM89, BLE87, BE86, BP87, Cac83, Cac87, CP89, CR86, CF83, CH89b, CD82, Cla89, Dan89, DdO89, DJ89, EN84, FF89, GR89, Hat86, HA87, HC89, KS89, Kha86, KM87a, Kor85, KS88, Kou85, KS87b, LS88, Lin88, Lin89, LE87, NN84, Oja83, PAT88, Phi89, SOY85, SYM86, Sha89b, Sta89, Swe83, SM85, TL89, Wil88, YM89]. divergence [Saw82]. divisibility [Dan89]. divisible [Dan88, Kag85, KS89, LE87, Vit88]. DMRL [AEN86, CH89b]. do [Sen87]. does [The87]. domain [Eub82, Tak87]. domains [Mee86]. Dominance [Nic88]. dominating [Li84]. dose [SKM86]. doubly [Kor87b]. drift [SB88]. durations [MS86]. Durbin [Ryc87]. dynamic [Bol89]. Dynamics [And84]. each [TC88]. early [Dal88]. Easily [Sim87b]. Easy [JR86]. Edgeworth [Ray89]. Editorial [Ano82a, Ano85b, Ano87b, Ano88b]. Effect [Ksh89, MG88]. effects [DGG83, Khu87]. efficiency [HR89a, Rot83, Rot84]. Efficient [Alb84, Bab86, KL85a, Wel88]. Efron [Vit88]. Eggenberger [Gou89b]. eigenfunctions [Ric83]. Eigenvalue [KPW84, Wan85]. Eigenvalue-eigenvector [KPW84]. eigenvector [KPW84, Wan85]. elementary [Sim83]. Elisa [GJ88a]. elliptical [KM87a]. elliptically [BB86]. EM-Algorithm [ZR86]. Embedding [Gou89a]. empirical [CH86, DT88, ER87, Ein87, Kuo86, Li85, Mar89, Wel84, Wel86]. entropy [CT82, Par82, RV85, The82]. epidemic [BGM^{+}86]. equal [Cha86]. equality [CS85, Hol87, NN84]. equally [Par88]. equation [FKT83, FRT84, Kou87, Nua87, TRF84]. equations [Dab89, HI83, Llo87, Oks87, SS86]. equireplicate [SJ88, Sin89]. equivalence [Sim83, Sim86]. Errata [Ano86a, Ano86b, Ano87c, Ano88c, Ano88d]. Erratum [Ano84b, Ano86c, Ano87d, Ano87e, Ano87f, Ano88e, Ano89c, Ano89d]. error [AB86, GJ88a, Kou85, Tsa89]. errors [AH89b, Fra88, JP84, Nay88, Rao84a, RB89, TRF84]. escape [Kor89]. estimate [CR89, Chi89, Dev89a, Dev89b, Fab88, Fra88, Geo88, GP85, RT86, Rou89b, She86]. estimated [HM87b, Koz85, SS84]. Estimates [Liu86, And84, Bur89, CH89a, CR89, Geo84, IR87, Koe84, Sin87, Wel84]. Estimating [Wei87, Yao88, Llo87, Nay88]. Estimation [BH86b, Dia87, FGS82, HM87a, JJ87, LB83, Alb84, Bab86, Bon86, CL87, CS88, CM88a, DdO89, Dey87, Fal86, Fer87, Feu88, FKT83, FFT84, FRT84, GMS87, GG89, GC86, Gup83, HH88, HW88, Hun85, Kea84, iKB87a, iKB87b, Kor87b, Kou85, Kub89, Kuo86, Le 88, Li85, Mar89, MN88, Mec85, Pet89, Rao84b, RE89, Rou89a, Rüs87, Sam89, Sch89b, Ses88, Sha89b, Sim87a, Sin86, TS89, WK83, Wei86, Wel88, Whi87, ZR86, dO89, dP85]. estimator [Bon89, CF86, Cla89, FK84, Fu89, KV85, Kni89, Leś89, Mie87b, Nic88, PT88, Rao84a, Rao85, Sta86, Sun85, Ver87, Yin89]. Estimators [Cha85b, KS83, Abd89, Ano89d, AK83, BB87, Ber89, Bra83, Cha85a, Cha86, CS89, CG84, EG86, Fu85, Gaj87, Har84b, JKV88, KL85a, Lar84, Li83, Li84, LP87, MM88a, Mas87, MM87, MS88, Mie87a, Pie87, Raf87, RB89, RR83, Ruk83, Sar86, SB88, Shi88, SS84, Ste86, VM83]. Evaluation [And89]. events [PX89]. Exact [Fra88, BM87, Khu87]. example [Eva89, LB83]. examples [KS89]. exchangeable [Dab88, Sca85]. Existence [BLE87, Kag88, AW84, Cha88]. exit [Mar83, Rus88]. exits [Rus88]. expansion [Kem88, Led87]. expansions [Ray89]. expectation [BE86]. expectations [Hoo89]. Expected [Pes82, And89, BB89, Bon89, MS86, Ric83]. Experimental [IO85, Kag88]. experiments [Lus89]. explanatory [Oks87]. exploration [BS84]. exponential [BLE87, CH89b, CS87b, Fu85, Gat88, Hou85, Kam82, KD85, KS88, LS88, Li87, MB87, MFP89, NN84, OR89, O’N85, PA88, RS88, Rao84a, Sam83, Sta86, Sta87b, Sta87a, Swe83, TL89, Yca89, YM89]. exponentiality [Mar83]. extended [HA87, Hun87]. extensible [Sca85]. extension [Kre88]. external [PS83]. extrapolation [HM87b]. extremal [Gou89a, SM85]. extreme [Bal87, DdO89, FGS82, GR89, Kea84, Wig85]. extreme-value [DdO89, Kea84]. Extremes [CD82, dO89].

factor [IO85, Iha87, Kan84]. factorial [Jon87a, Kem88]. failure [CKS88, GM85, JC89, MS85, MM87]. failure-rate [MM87]. fallacy [Aka82]. families [CS87b, Hou85, MFP89, RS88, Yca89]. family [Ath86, GR89, KS87b, LS88, Li83, Li84]. favorable [Sim87b]. feasible [Kor87a]. fewer [Kag88]. fields [Bra89]. filter [RB87]. filtering [CKS88]. filtrations [MW87]. financial [AC82]. finite [Dan89, Eve88, Gup83, Hun86, Ric85, Rod87, RE89, She86, Yca89, ZR86]. finite-dimensional [Ric85]. Finitely [Str89, RS86b]. Finiteness [GW89]. First [Pet89, Har84a, Le 88, Ryc87]. first-order [Le 88]. first-passage [Ryc87]. Fisher [Riv86]. fit [BG88b, CS87b, Ray88, Riv86, Sve85]. five [Fin87]. fixed [Geo88, Rou89a]. flow [KL85b]. Fluctuation [Edl84, Ros83, Ros84]. fold [Khu87, Tsu88a]. footrule [SS83, Fra88]. forecast [Tho89]. forecasting [JP84]. form [AH89a, Ein87, KL85a]. formation [AEN86]. forms [Jon87b]. formula [Bru88, Ryc87]. forward [Tay82]. four [Tsu88a]. four-fold [Tsu88a]. framework [YR88]. free [CD88, RE89, Swe83]. frequencies [Bon89]. Friedman [Rot83]. function [Ath87, Bab86, BP87, CL87, Dia87, Eva89, Fal86, Gaj87, Geo88, Kor87b, Krz83, Leś89, Liu86, MM87, RB89, Rou89b, Sca85, Sin86, Vit88, Wel84, Wel86, Wel88]. Functional [Sha89a, Dab85, DT88]. functionals [Par83b, Rüs87, Smo85]. functions [Bar87, BH86b, CP85, CP86, Dev84, Gaj87, Geo84, Mai88, MM88b, Mis89, Ric85, Ver88, Wei87]. fundamental [CF86]. Further [SJ88].

G [Ano86d]. gain [Wei87]. Galambos [TX89]. game [GH84, San88]. gamma [AB86, And84, CR86, HC89]. Gaussian [FB86, Gzy87, Kle86, LM87, Lus89, LW82, Rao84a, Ses88, Sin86]. general [BGM^{+}86, Cha88, Ein87, Lar84]. Generalization [Dan88, Pap88]. generalized [AB86, Cac87, Fam87, Gou89b, GH84, Har89, Ksh89, PX86, PGP83, Ver87]. generating [Dev84]. generator [Cha89, Zie86]. geometric [Arn88, PGP83, Sen87, Sze88]. Georg [Ano89d]. given [MS86]. Godambe [Bha87, God87a, Jos87]. Gompertz [Tan86]. goodness [BG88b, CS87b, Ray88, Riv86, Sve85]. goodness-of-fit [BG88b, Sve85]. Graphical [Tsu88a, Sri84]. green [CL87]. group [FM87, Kag85, PS83, Yao89]. grouped [GJ88b, HR89a]. grouping [Eub82, tHK84]. groupings [LC82]. growth [God87b, Hüs85]. Gumbel [GR89, Tak87].

Hadamard [Sin88]. Haight [LS88, SH89]. Hájek [Par83a]. Half [Sha89b]. Half-sample [Sha89b]. Hans [Ano89d]. Hans-Georg [Ano89d]. harmonic [Sen87]. Harris [Tay82]. having [Pet89]. hazard [Bar87]. hazards [CL87, CM88a, GM85, O’n86]. head [MS84]. heat [KL85b]. Hermite [GP85]. heteroscedastic [Tsa89]. Hierarchical [Per82]. histogram [KV85, Lec85]. histograms [RV85]. historical [SKM86]. Holdings [AC82]. homogeneity [BB87, BH86a, CD88, GJ88a]. homogeneous [DM84]. Hotelling [Ksh89]. hypergeometric [BP87]. hypotheses [Saw82]. Hypothesis [Che84, BH86a, HR87].

i.i.d [GW89, Ros84]. I.I.D. [Ros83]. identically [Hor85]. identifiability [LE87]. identification [AB89, Dab89]. identifying [Gho88]. II [Ric85]. Implementing [Wel86]. implications [Kre88, LW82]. implies [CKZ86]. importance [Bue88]. improper [Iha87]. Improved [Dey87, Kub89, Sch87, SYM86]. improvements [YM84]. Improving [PT86, Gaj87, MM87]. Imputation [TS89]. income [Kor85]. incomes [Kor85]. Incomplete [MM84, Shi89]. Inconsistencies [Bha89]. Inconsistency [O’n86]. increment [Rhe86]. increments [PA88]. independence [Baj87, BH84, BH86a, CKZ86, Feu88, HR89b, MS84, O’B82, O’N85, Rud84, Sam83]. Independent [Swe83, Hor85, MS88, PA88, RW85, TG84, The87]. index [Ano83, Ano84a, Ano85a, Ano87a, Ano88a, Ano89a, Ano89b, BP85, Led87, dP87]. indexed [GL89]. individual [HR89a]. ineffectiveness [Pic84]. inequalities [sL89, RT86, RS89, TX89]. inequality [Fu83, KS87b, Pap88, Vit88, Wit85, YSM87]. inference [AH89b, KL85b, MS85, Pic84, TH87]. Infinite [AB89, Dan89, GK89, HH88, Lus89, Zie86]. infinite-dimensional [HH88]. infinitely [Dan88, KS89, LE87, Vit88]. Influence [Kao85, Shi89, BH86b, Par83a]. influential [Gho88]. information [FFT84, GG89, Gup83, SKM86, The87, TC88, Whi87]. Inheriting [BH84]. insensitivity [MS87a]. inspection [Kre88]. integrability [CS87a]. integral [MB87]. integrated [HM87a]. intensity [Leś89, RB89]. intermediate [Coo88]. internal [PS83]. interpolated [HS86]. interpretation [FK84, Hat86, Par82]. interrelationships [Tay82]. interval [DLS86, SM87]. intervals [BB88, BB89, HM89, Her88, HS86, Ray89, TBGG89]. Intrinsic [dO89]. invariance [Arn85, Dab88, Par83b]. Inverse [Jon87a, CT82, FB86, Ses88, Sin86, Sun85, WD88]. inversion [Ray89]. iterated [Dab85, DT88, Deh89, Ein87]. Iterating [Hop85].

jackknife [Shi88, RT86]. James [Nic88]. joint [Kor85]. jump [CGN83].

Kalman [RB87]. Kaplan [Mie87b]. kappa [Agr89]. Kendall [NY88]. Kent [Riv86]. Kernel [Geo84, Abd89, Bra83, Chi89, Dev89a, Dev89b, HW88, Har84b, Leś89, LP87, MM87, Mül84, OT87, Sch89b, Sin87, Ver87]. kernel-type [Bra83, LP87, Ver87]. kernels [MN88]. Kersting [Pfl83]. Kiefer [Lev84]. Kim [Fu89]. Klass [God87b]. known [Hol87]. Knut [Afs86, Afs87, Afs88b]. Kolmogorov [KS85, Sen84a, YSM87]. koziol [CL87]. koziol-green [CL87]. Kronecker [DPV84]. kurtoses [BB87]. kurtosis [BB87, RS89, Sri84].

Large [Bir88, HH88, HR87, Kok87, LM87, RS86b, Wit85, CHR86, Fu83, JKV88, Led87, Par85, RR83, Sch89b]. last [Rus88]. lattice [WJ88]. Law [Bir88, HR87, Kok87, LM87, RS86b, Wit85, Ath83, Dab85, DT88, Deh89, Ein87, Wes87]. leading [Bha89]. least [DGG83, EG86, Hun87, IO85, MS87b, OI83, Rao84a, Rao85, dP84]. least-squares [EG86, IO85, OI83]. length [Pap84, Sen87]. length-biased [Sen87]. Letters [Ano89d]. Levene [CS85]. Lévy [Sim86]. ‘Lévy’s [Sim83]. life [Abo88, AM89, AK83, EN84]. Likelihood [SR84, Zel88, Aka82, BP87, CF86, CL87, CJT87, Eva89, FK84, Fu85, Fu89, Hol87, IO85, Iha87, Kou87, Le 88, Pic84, Sar86, WK83, Zel86]. Limit [AJZ89, Joh88, Pfl83, Rhe84, Rhe86, iYT87, AA88, DJ89, DLS86, JR86, KS89, Mut88, Yin89]. limiting [JJ84, Sen84b]. limits [CM84, Har89]. line [Rao84b, RP85, Yin89]. Linear [Mar89, dP84, AB86, AH89b, Bon85, BBW89, Bon89, Cha85a, Dab89, Koe84, KS83, KS85, Kou85, LM87, Led87, Nua87, Pie87, Rot84, Shi89, Smo85, dP85]. linearity [Csö85]. Lipschitz [SS86]. Local [Baj87, CHR86, CH89c, Fab88, HS89, SS86]. Locally [Rud84, Sch89b]. location [Che84, Fu85, GC86, HR89a, MS88]. log [AB86, BBW89, Bon89, Zel86]. log-likelihood [Zel86]. log-linear [AB86, BBW89, Bon89]. logarithm [Dab85, DT88, Deh89, Ein87]. logarithmic [Fam87]. logistic [Arn88]. lognormal [Kap88]. long [DT88]. long-range [DT88]. longest [PM86a, PM86b]. look [LY89]. Lorenz [AV86]. loss [And84, Hsi89, Kub89, MM88b]. low [Mut88]. Lower [Sca85, Dev89b, JH84, MM88b, Rot83]. lumber [JH84].

Mack [Ano89d]. magic [Hed87]. main [MG88]. maintenance [CKS88]. many [RS86b, The87]. mapping [Mut88]. marginal [CD82, Har84a]. market [GL89]. market-indexed [GL89]. Markov [AT89, AC82, Ath85, Bur89, CGN83, DM84, Gzy87, Joh88, KS89, SS83, Tsa85, Yca89]. Marquardt [OI83]. Martingale [RT86, EK88, Gou89b, Llo87]. martingales [LM87]. match [Hol89, RT83]. matche [JJ84]. matched [Tsa85]. matching [Pes84a]. mathematical [Per82]. matrices [Flu86b, Flu86a, Hög87, Hol87, KPW84, Sin88, Wan85]. matrix [BH84, CT82, Dey87, Eub84, Kni89, Kub89, Phi89, Sen84b, WD88]. Maxima [HR89b, God87b, Hüs85, Roo84]. Maximum [CL87, Par82, RV85, SOY85, BP87, CF86, CT82, FK84, Fu89, Hsi89, IO85, Iha87, Le 88, Sar86, The82, WK83, Wei86]. McCarty [YM84]. Mean [RB89, Abo88, AK83, BB89, CR86, CS89, FGS82, GMS87, Li83, Li84, Li85, Sta86, Sta87a, Wu89]. means [AV86, CM88b, Hol87, Hun86, Kuo86, LW82, Sen87, Wei87]. measurable [Smo85]. measure [Kle86, OT87, Sri84]. measurements [Afs88a]. measures [Bök88, GG89, Pic84]. measuring [And86]. Median [CG84, CR86, She86, SM87]. medians [AV86, LH89]. Meier [Mie87b]. memory [Hsi89]. mesure [CHR86]. Method [Fu89, Ald89, HM89, Kap88, Kor87a, KM87b, Sri84]. Methodologies [Fer87]. Methods [Dev84, Feu88, Fin87, Gho88, Hun87, IO85, SM87, iYT87]. metric [Str89]. Michael [Ano86d]. minima [Gou89a]. minimal [God87b]. Minimax [HM87b, Fab88, Li83, Li84, MS88, Ruk83]. minimization [Arn88, HW88]. minimizing [BH86a]. Minimum [FFT84, Jey85, Kou85, Cla89, Mec85]. misclassification [CR89]. Mise [Sin87]. Mises [Pap84, Rüs87]. missing [Fer87, How83, Ksh89, TS89, ZR86]. misspecified [O’n86]. Mixed [Bon86, Eve88]. mixed-mode [Eve88]. mixing [AP86, Bra89, DJ89, IR87, Mas87]. mixture [BLE87, Cla89, Eve88, Hat86]. mixtures [Cac83, LE87, MS88, Zel86, Zel88]. MLE [MMR88]. modal [Har84b]. mode [Eve88]. model [AB86, Agr83, Agr89, AA89, Ber89, Bon85, BBW89, CL87, CKS88, CM88a, Eve88, FB86, Gou89b, JC89, Kan84, Khu87, Kor85, KM87b, LR89, Leś89, O’n86, Pet86, Pie87, Rao84a, Rao85, RP85, RB87, Rod87, Ron85, Ser84, Tsa89, Wal85, ZR86]. modelling [SKM86]. models [AH89b, Bec89, BM87, Bol89, Bon86, Bon89, Cha85b, Feu88, FFT84, GM85, HH88, Kao85, Kap88, Koe84, NH89, Ric86, Rüs89, Sve85, Tsa85, Tsu88b, WK83, Web84]. Moderate [JKV88]. Modified [Ath87, Riv86]. modulus [ER87]. Moivre [JR86]. Moment [sL89, Wit85]. Moments [BB86, Abd89, BLE87, Dal87, GW89, Jon87a, Jon87b]. monitoring [CZ84]. monotone [Bue88, MS85, Wef85]. monotonic [Gaj87]. Monro [Gol88]. Monte [Fin87]. motion [CHR86, Sim83]. moving [Rus88]. MSE [Sun85]. MSE-superior [Sun85]. Müller [Ano89d]. multi [PT86]. multi-way [PT86]. Multicollinear [DGG83]. multidimensional [Dav87]. multinomial [Ahm85, BP87, Bon85, iKB87a, iKB87b, Zel86]. multinormal [Dey87, Rod87, SR84, ZR86]. multiple [Li87, PX89, Saw82, Shi89, The87]. multiplicative [Leś89]. multistate [Bue88]. Multivariate [BE86, PAT88, BP87, Cac87, ER87, Ein87, EN84, FF89, Geo88, Gup82, Hol87, Li83, Li84, OR89, Oja83, RS86a, Sam84, Sri84, SH87, Sun85, Sve85, Wig85]. must [CHR86]. mutually [Afs87].

Nadaraya [Geo88]. Narrow [Ver88]. natural [LS88, Tsu88b]. NBUE [AEN86]. near [GJ88b, JJ84]. near-matche [JJ84]. Nearest [Cox83, CH89a, Mie87a]. necessary [Bal89]. negative [BSS85, Bon85, Lin89]. neighbor [CH89a, Mie87a]. neighbours [Cox83]. nested [BM87, Khu87]. nesting [Khu87]. network [CZ84]. Neyman [Akr88]. nodes [Mut88]. noise [Chi89]. Non [LW82, Sam89, BM87, DdO89, Dev89a, Yao89]. non-consistency [Dev89a]. Non-Gaussian [LW82]. non-nested [BM87]. Non-parametric [Sam89, DdO89]. non-randomized [Yao89]. non-zero [LW82]. noncentral [BG88a]. nonhomogeneous [RB89]. noninfinite [Dan89]. nonlinear [Feu88, OI83, Rao84a, Rao85, Tsa89]. nonnegative [MM87, Sen84b, TBGG89]. nonnested [AT85]. Nonparametric [AJ89, JC89, Kor87b, Raf87, AT85, Ano89d, HR89a, KM87b, MM88a, Mül84, PT88, Ray89, Rot84]. nonregular [Cha85a, Rao85]. Nonstationary [HI83]. norm [Akr82, FF89, Mis89]. normal [Ath86, Bec89, Bol89, CM84, Cla89, Fra88, GMS87, Hol87, tHK84, HR89b, Jon87b, Li83, Li84, MS88, RS86a, Ruk83, TRF84, Wei87, Wes87]. normality [BR88, Bra83, GJ88b, HK88, KM87a, KV85, Sri84, SH87, Wel84]. Normalizing [Tak87]. Note [HC89, IR87, KS88, Sta87b, Akr82, AA88, AP86, Bar87, Bir88, BG88a, CS85, Dab82, Fam87, FK84, HM89, Hög87, Hol89, tHK84, Hun87, JR86, Jos87, Kea84, Kem88, Ket89, Kni89, Koc89, Koe84, Koz85, Kuo86, Le 88, Leś89, Nay88, Oks87, Par83a, Pet86, RL83, Ryc87, SYM86, Shi88, Sze88, WK83, Wei86, Wel84, Whi87, Yin89, YSM87, ZR86]. nuisance [PT86]. number [AA88, Cha89, Fin87, JJ84, LB83, Mut88, Nay88, Rus88, Yao88, Zie86]. numbers [Gup82, Bir88, HR87, Kok87, LM87, RS86b, Wit85].

observation [FGS82, Ksh89]. observations [Ber83, Fer87, Gho88, Sar86, Shi89, Wu89]. observed [LB83]. obtain [CF83]. obtaining [Wig85]. Occupation [MS87c]. occurrence [Dal88]. occurrences [PX89]. OLS [Bal89]. One [RT83, TBGG89, TH87]. One- [RT83]. one-sample [TH87]. One-sided [TBGG89]. operators [Ric83]. optima [Pet89]. Optimal [CZ84, FB86, Lar84, Llo87, Mül84, Tsu88b, AW84, Cha89, Eub82, Lec85, Raf87, Rud84, Sch89b]. optimality [Baj87, CF86, Hed87, MG88, Ste88, Tal86]. Optimum [Kat89, tHK84, SS84]. order [Che84, Coo88, Gat88, Har84a, HS86, HA87, Hoo89, HM87b, KL85a, Le 88, Lin88, Lin89, Pet89, PAT88, Rei86, dP87]. ordered [CD88, Lee88, Sim87a]. Ordering [AAO89, AK88, AV86, Bar86, Bar87, BSS85, PS83, Pes82, Pes84b, Tsu88b]. orderings [BH84]. ordinal [Bec89]. orthogonal [Afs86, Afs87, CP86, Raf87]. oscillation [ER87]. other [Wil88]. outcome [MS86]. overlapping [Ray88].

P [Ano89d]. pairs [Tsa85]. pairwise [Pes84b, RW85]. parabolic [AB89]. paradox [Bha87, Jos87, Kre88, God87a]. parallel [AEN86]. Parameter [Dab89, Agr83, Agr89, AB89, Cla89, JH84, NN84, SB88, Sto83]. parameters [Che84, Fu85, KS83, Koz85, PT86, Ser84]. parametric [DdO89, Sam89]. Pareto [Arn87, AH89a]. part [Nic88]. Partial [Dav87, DPV84, Ebe84, God87b, Hüs85, NH89, NY88, Rus88, SR84]. partially [Hol87, Ric86]. parts [JP88]. Pascal [Wil88]. passage [Ryc87]. path [KS87a]. patterned [KPW84, Wan85]. Paul [Ano82b]. Pearson [Bon89]. penalized [EG86]. percentile [HM89]. percentiles [Kea84]. perfect [Dav86]. period [Zie86]. periodic [Leś89]. periods [SH89]. perspective [Eub82]. phase [OR89]. phenomenon [RL83]. Philipp [Dab82]. Pitman [Rot83]. Plane [Iac85, Are89, SS86]. play [MS86]. players [San88]. plot [MM84]. point [Are89, Bab86, Cac83, Cac87, Cha86, RR83, Wel88]. points [LB83, Rou89a, Yao88]. Pointwise [Bur85, GP85]. Poisson [Ros84, Ahm85, And84, BBW89, CR86, Dal88, Dia87, Edl84, Har89, Li85, RB89, Ros83, Sti82, Wil88]. Pólya [AB88, Dev84, Gou89b]. polynomials [CP86]. Pool [DG88]. Pool-Adjacent-Violator [DG88]. Population [Bol89, Gup83, Hun86, JP84, Pap84, RE89, VM83, Wu89]. populations [CC89, Hol87, RP85, Rod87, ZR86]. Positive [Mis89, Ric85, KD85, Nic88]. positive-part [Nic88]. positively [Bir88]. possible [CH89c]. posterior [MS87a]. power [tHK84, LE87]. Practical [LW82]. precision [Dey87]. prediction [Bol89, Hun86, LC82, ZR86]. predictive [CJT87]. predictors [Rod87]. Preface [Joh82]. presence [FGS82]. preservation [Sze88]. pricing [GL89]. principal [MG85, Sch87, Sch89a]. principle [Aka82, Bha87, Par83b, ZR86]. principles [Dab88]. prior [MS87a]. priority [Ros84]. priors [And84]. privacy [Alb84]. probabilities [Ahm85, Ber83, GS87, RR83, YR88]. Probability [Ano89d, Sim87a, Abd89, AAO89, Ath87, BS84, Bra83, Bur85, CR89, Fu83, MS87c, PX89, Wei86]. problem [Alb84, Ald89, Hol89, KL85b, MS87c, OI83]. problems [Eub82, Har89, MMR88]. procedure [CS88, DM84, Gol88, Hus88, Lev84]. procedures [Pes82, Pes84b, Wel86, Yao89]. process [Arn88, AH89a, DT88, Dia87, ER87, Ein87, Har84a, Le 88, Llo87, PA88, RB89, RW85, Sca85, Sta87b, Tan86]. processes [And84, Are89, AH89a, AP86, Cha85a, CD82, Coo88, CGN83, CH86, DM84, Gou89a, Gzy87, KS89, Kuo86, Mas87, Ste88, Str89, Tay82, Yan89, Yca89, dP85, dP87]. product [HW85, HW86, Lee88, TG84, Yin89]. product-limit [Yin89]. products [Hög87, Sen84b]. projections [Par83a]. proof [Akr84, AT89, Ano86d, EK88, GK89, Joh88, Pfl83, Pic84, Rei86]. proofs [JR86]. Properties [HA87, Bur85, Gat88, Geo88, Hus88, NY88, PGP83, Rou89b, Sar86, Yin89]. property [Fab88, Hsi89, JH84, Kam82, Pla88, RS88, Rot84, SS83, SM85]. proportion [Cla89, Kor85]. proportional [CL87, CM88a, GM85, JJ87, O’n86]. Proportionality [Flu86a]. provided [TC88]. proximity [CH89b]. pure [MS84].

quadratic [Jon87b, Kub89, Ste86]. qualitative [tHK84]. quantile [Bab86, CH86, Eub82, Fal86, HM89, LP87, PT88, Ruk83, RC87]. quantiles [AJ89, Mar89, SAU84, Sam89, Ver87]. quantizers [AW84]. quasi [Agr83, Zie86]. quasi-random [Zie86]. quasi-symmetry [Agr83]. question [FS85]. questions [LC82, Pes82, Pes84b]. queues [Li87]. quotients [Phi89].

Radial [JP88]. radius [LB83]. Random [How83, Alb88, AJZ89, BB86, Bir88, Bra89, CP85, CP86, Cha89, Dan88, Dev84, FS85, FM87, GK89, GW89, Gri88, HW85, HW86, Har89, Hög87, Hor85, HR89b, IR87, JP88, Khu87, Len87, Liu86, Mec85, MS86, MS87c, Mut88, Pap84, Ros83, Ros84, Rüs89, RC87, Sta86, Sta89, TG84, WD88, Zie86]. randomized [Mat84, Yao89]. randomly [GW89]. range [Ath85, DT88]. Rank [Alb88, AH89b, Tsa85, Baj87, Bru88, Led87, NY88, Rud84, TH87]. Rank-based [AH89b]. rankings [PS83]. Rao [MM88b]. Rate [Bur89, And84, FK84, God87b, Hüs85, MM87, Raf87, Rao84a, Rhe84]. Rates [Kok87, Fu85, GJ88a, MS85, MM87, Ric86, Roo84, RR83, Ste88]. ratio [BB89, Fu85, Hol87, Sam84, SR84, Zel86, Zel88]. ratios [Jon87b]. recapture [Nay88]. reciprocal [Bab86]. Rectangular [Sin88, Abd85, Sta89, SM85]. recurrence [Bru88, Tay82]. recurrent [Ath85]. recursive [Fab88, Mas87]. reduce [CG84]. redundancy [Flu86b, Kha86]. region [HM87b]. regions [Jey85]. Regression [Hun85, Alb88, Ano89d, AJ89, Cac83, Cac87, CH89a, Cha88, Dia87, DGG83, FM87, FB86, Geo88, Kao85, KS83, KS85, Kou85, LM87, MM88a, MG85, Mül84, Oks87, Raf87, Rao84a, Rao85, Ron85, Rou89a, Shi89, Sun85, The87, Tsa89, Wal85, Web84, dP85]. regressional [Wes87]. regressions [AT85]. Regular [Mee86, Ebe84, MFP89]. related [RC87]. relation [CK87]. Relations [TC88]. Relationship [Dav86, GM85]. relationships [Eub82, SKM86]. relative [tHK84, PS83]. reliability [BE83, Kea84, Mai88, Par88, Sin86]. reliable [Par88]. remaining [Abo88]. remark [Mie87b, Ros83, Ros84, Wil88]. Remarks [BS84, AW84, Bök88, Nua87, SS86]. renewal [Kle86, Sta87b, Ste88, Yan89]. renewals [AA88, Kha86]. repeated [Afs88a, RB87]. replications [Fin87]. reported [Kor85]. representation [EK88]. Representations [Ric83, Len87]. reproductive [RS88]. resampling [Web84]. residual [Oks87]. residuals [Tsu88a]. Resolution [God87a, Har84b, Jos87]. response [Mat84, Pet89, SKM86]. responses [HM87b]. restored [How83]. restricted [GMS87, dP85]. restrictions [Alb84, dP84, dP87]. result [BLE87, EK88]. results [BBW89, LP87, The82, Wal85]. ridge [Cha88]. right [GC86, PT88, Sta86]. right-censored [GC86, PT88]. right-censoring [Sta86]. risks [YR88]. Robbins [Gol88]. Robust [Ron85, GC86]. Robustness [Pes84b, SGS87, Par85]. role [BS84]. Rosenthal [Wit85]. roulette [GH84]. row [DPV84, SGS87]. row-column [SGS87]. ruin [MS87c]. rule [Bha89]. rules [CR89]. runs [PM86a, PM86b].

Sabatier [Ano82b]. Saddlepoint [Hou85]. same [Coo88]. sample [Abd89, Alb88, And86, AJZ89, CM84, Dal87, LR89, LH89, Par85, RT83, SAU84, Sha89b, She86, SS84, TH87, Wei86]. samples [FKT83, Hoo89, Rao84b, Riv86, Sch89b]. Sampling [Cha85a, BP85, GG89, Hun87, Iac85, Nay88, RP85, RE89, Sen87, Sha89b, Tal86, ZR86]. Sanov [Fu83]. Savage [Akr84, Ano86d]. scale [Che84, KS83, Swe83, Wel84]. scale-free [Swe83]. scaling [Tsu88b]. scheme [DLS86]. Schur [BP87]. Schwarz [Pap88, Yao88]. score [Liu86]. scores [KD85, Tsu88b]. screening [KM87b]. seasonal [LR89]. Second [dP87, HM87b, Pet89]. second-order [HM87b]. seemingly [Oks87]. selected [CS89, SAU84]. Selecting [MG85]. selection [Chi89, Lev84, Pes82, Pet86, Ron85]. Self [Afs86, Sze88]. self-decomposability [Sze88]. Semi [Afs88b, DM84, Tsa85]. semi-Markov [DM84, Tsa85]. sensitivity [GJ88a]. separate [Saw82]. Separating [OT87]. sequences [CGN83, Dab82, Dab85, Dab88, DT88, DJ89, Ebe84, How83, Kle86, iYT87]. Sequential [Mec85, Hus88, Lev84, MMR88]. series [Cha85b, Fam87, Fer87, Feu88, God87b, GP85, LW82, LE87, Raf87]. servers [Li87]. set [Kor89, Pie87, Sen84b, TC88, Ver88, Yca89, dP87]. sets [Gho88, TC88]. setting [CM84]. settings [Mat84]. several [AT85, Hol87]. shape [Gup82]. Shapiro [SH87]. Sharp [RS89]. shift [Lus89]. short [EK88, Fam87, Joh88]. shrunken [Rod87]. sided [TBGG89]. sigma [MW87]. signed [TH87]. significance [SAU84]. Simple [AK83, Agr83, Akr84, Ano86d, CM88a, GK89, GG89, Kap88, OI83, Pie87, Yao89]. simple-stage [GG89]. Simplicity [Bro82]. simulation [Fin87]. Simultaneous [FKT83]. simultaneously [AT85, Gup82]. Singh [WJ88]. single [RB87]. singular [Pic84]. situations [Sun85]. size [AJZ89, Gol88]. sizes [CM84, FFT84, SJ88]. skewness [RS89, Sri84]. small [GJ88a, Riv86, Wei86]. small-sample [Wei86]. Smirnov [KS85, Sen84a]. smooth [PT88]. smoothing [Eub84, Leś89, Sim87a, Yan88]. Sobel [Lev84]. solution [BGM^{+}86, Kou87]. solutions [Iha87]. Some [AW84, BT84, Bök88, Lee88, LP87, NY88, Nua87, Par85, Rod87, Rou89b, SS86, Sin89, Sta89, TX89, The82, YM84, Baj87, Ber89, Bra83, DT88, GG89, KS89, Pet86, PGP83, Wil88, iYT87]. space [PA88, Rhe84]. spaces [Ric85, Str89]. spacing [Eub82]. spacings [AJZ89, Swe83]. sparse [Rei86, Sim87a, Zel86]. sparsity [Wel88]. Spearman [Fra88, SS83]. special [Dan89]. Spectral [Len87, Par82, WK83]. Spherical [Phi89]. spline [Bur89]. splined [Ric86]. splines [Eub84, Yan88]. split [MM84]. Splitting [Her88, DLS86]. spring [Kat89]. SPRT [Sta87a]. square [BG88a, BH86a, Bon89]. squared [HM87a, RB89, Sch87]. squaredness [BH84]. squares [BB89, DGG83, EG86, Hed87, Hun87, IO85, OI83, Rao84a, Rao85, dP84]. SR [Kag85]. Stable [DJ89]. stage [CS89, GG89, RB87, Ser84]. standard [Sun85, TRF84]. standardized [Dal87]. stationary [AH89a, CGN83, Gzy87, Mas87, RW85]. statistic [And86, Bru88, Pes84a, Ses88, SH87, dR88]. statistical [Ber89, Hed87, KL85b, LM87, Par83b, Sha89a, Sto83, YR88]. Statistics [Ano89d, AJZ89, Che84, Deh89, Gat88, HS86, Hoo89, KL85a, KD85, Kok87, Led87, Lee88, Oja83, RT83, Rei86, Riv86, TH87]. Stein [Ald89, Nic88, Vit88]. step [Gol88, dP84]. Stirling [Kem88]. stochastic [AB89, AH89a, BE83, BSS85, Coo88, Dab89, Ebr85, Kor87a, KL85b, Llo87, Nua87, PA88, RW85, SS86, Ser84, Tan86, dP85]. stopped [GW89, Hor84]. stopping [MS84]. Strassen [Dab88]. Strassen-type [Dab88]. strata [Mut88]. stratification [Eub82]. stratified [Tsu88a]. strength [AA89, Ebr85]. stress [AA89, Ebr85]. stress-strength [AA89, Ebr85]. strictly [Dav86]. Strong [Abd89, Ath83, Hor84, MMR88, AP86, Gou89b, Kou87, Mie87b, Ste88, Bir88, HR87, Kok87, LM87, RS86b, Wit85]. strongly [DJ89, RS88]. structural [Ber89, Kag85]. structure [Bon86, Iha87, Tsa89, dO89]. structures [CKZ86, Per82]. study [Fin87]. stuttering [PX86]. subject [Bon89]. subsequences [Hüs85]. subspace [Sch89a]. Successes [PM86a, PM86b]. succession [Bha89]. Sufficiency [Arn85, Rei86]. sufficient [Bal89, Dan88]. suggestions [Ebr85]. sum [Bru88, DPV84]. summability [iYT87]. sums [Ebe84, GW89, Gou89a, HK88, Hor84, Hor85, NH89, Rus88, Sta89, TBGG89]. superior [Sun85]. superpopulation [Bol89, Rod87, ZR86]. sure [ER87, Mas87]. surface [Pet89]. survey [LC82]. surveys [RB87, SS84]. survival [Ber83, CL87, GM85, Rou89b, YR88]. survivorship [Kor87b]. symmetric [FF89, FS85, HW85, HW86, Ric85, Wei87]. Symmetrized [CH89a]. symmetry [Agr83, Koz85, O’N85]. system [Ebr85, Nay88]. systems [AEN86, AB89, BE83, Bue88, FRT84, Par88, TRF84].

table [BH86a]. tables [Bök88, Dav86, Dav87, Fra88, Lee88, Sim87a, Tsu88a, Wal85]. tail [AAO89]. tails [CM84]. Tailweight [AM89]. techniques [Web84]. tell [Sen87]. test [Akr88, Bar86, BR88, Bon89, CC89, Flu86b, Hol87, Khu87, KS85, Lee88, NN84, Ray88, Sch87, Sch89a, Sen84a, Sto83, Sve85, Wel84, Wu89]. Testing [AT85, AK88, Csö85, GJ88a, O’N85, CD88, Che84, Feu88, KD85, Koz85, Riv86, Saw82, SM87, SR84, Yao89]. Tests [Gup82, SAU84, Alb88, AK83, Baj87, BM87, BW89, BB87, CS85, CD88, CS87b, GJ88a, HR89a, Rot83, Rot84, Rud84, Sam83, Tsa85, Zel88]. their [Geo84, Gup82, NH89, Tay82, Wel84, YR88]. theorem [Akr84, AT89, AA88, Ano86d, AB88, Cha88, Dab82, WJ88, Joh88, Pfl83, Rhe84, Rhe86, iYT87]. theorems [JR86, JKV88]. theory [BW89, Par85, Rao85, Sha89a]. three [JH84, San88]. three-parameter [JH84]. time [Bar86, Cha85b, CHR86, CH89c, Fer87, Feu88, GM85, GH84, JC89, Li87, LW82, RS86b, Rus88, Ser84]. time-dependent [Ser84]. times [Mar83, MS84, Tay82]. tolerance [CM84, JH84]. topologies [Whi87]. total [Bar86, VM83]. Toulouse [Ano82b]. trails [Sto83]. transect [Rao84b, RP85]. transformation [Bar86, Tho89]. transformations [dR88]. transformed [Hun85]. Transforming [GJ88b]. translation [Rüs89]. tree [AT89]. trend [Rot84]. trends [Lee88]. triangular [Deh89, Edl84, Joh85]. trimmed [HK88]. trimming [VM83]. trivariate [Sam83, Sim86]. true [Kor85]. truncated [Sta87b]. Two [CS89, Ebr85, Ser84, TH87, Ald89, BH86a, Cla89, EN84, Flu86b, Gho88, Khu87, NN84, RT83, RB87, TC88, dP84]. two-dimensional [Ald89]. two-fold [Khu87]. two-parameter [NN84]. Two-sample [TH87, RT83]. Two-stage [Ser84, RB87]. two-way [BH86a]. type [Ano89d, Bal87, Bra83, Dab88, HR87, JKV88, KS85, KS87b, LP87, MM88a, MM88b, OR89, Pes82, Pes84b, Rhe84, Rot83, Sen84a, TX89, Ver87].

ultimate [MS86]. unbalanced [Khu87]. Unbiased [Gup83, Rüs87, Cla89, CS89, Hol87, iKB87a, iKB87b]. uncorrelatedness [CKZ86]. under-reported [Kor85]. underadjustment [RL83]. underlying [Sto83]. undersized [FKT83]. unequal [Ber83, SJ88]. unidimensionality [Sto83]. Uniform [CM88b, Abd89, Afs88a, CD82, IR87, Mie87b, Swe83, TL89]. uniformly [WD88]. unimodal [SOY85, SYM86]. unimodality [BT84, Wef85]. unit [Pap84]. units [Kag88]. Universal [HK88, MG88, Dev89b]. Université [Ano82b]. unknown [Kou85]. unrelated [Oks87]. upon [KL85a]. upper [CP85, Fu83]. urn [Gou89b]. urns [Sim87b]. usable [Ray88]. Use [Mat84, LC82, SKM86, Wig85]. used [BE86]. useful [Tho89]. using [Arn88, AB88, BSS85, Che84, Gup83, SAU84, TS89, VM83].

vague [Ver88]. validated [Dev89a]. validation [HS89]. validity [DG88]. value [And89, Bal87, DdO89, GR89, Kea84, Ric83, ZR86]. valued [Gzy87, Le 88]. values [CF83, How83, SS84, TS89, Wig85]. variable [Dan88, Mec85, The87, Vit88]. variables [Bir88, CP85, CP86, FS85, Flu86b, GK89, GW89, HW85, HW86, Hor85, IR87, Jon87b, Kan84, Oks87, Ros83, Ros84, RC87, Sta87b, Sta89, TG84]. Variance [AB88, Pap88, Bol89, CP85, CP86, CP89, CS88, Gup83, JP84, Kag88, Khu87, MS88, RT86, SOY85, SYM86, She86, Shi88, TBGG89, Vit88]. variance-balanced [Kag88]. Variances [LH89, CS85]. variate [TC88]. variates [BB86, Dev84, tHK84, Hun85, MS88, TC88]. variation [AA88, Mee86]. various [Ste86]. vector [Gzy87, JP88, LM87]. vectors [HR89b, Oks87, Pap84]. version [Nic88, Sim86, Vit88]. via [Ath87, KS87b, Sim87a, Yao88]. Vik [Afs86, Afs87, Afs88b]. Violator [DG88]. voisinage [CHR86]. volume [Ano83, Ano84a, Ano85a, Ano88a, Ano89a, Ano89b, Jey85]. volumes [Ano87a].

waiting [GH84]. walk [Gri88, MS86, MS87c]. Walker [HI83, Kni89]. Waring [PX86]. Watson [Geo88]. way [BH86a, PT86]. Weak [AA89, Ebe84, AA88]. weakly [iYT87]. Weibull [CC89, JH84, Kap88]. weighing [CK87, Kat89]. weighted [CH86, DGG83, Ein87, MM88b]. weights [Lar84]. which [CKZ86, Gaj87, How83, Sim87b, Tak87]. whole [Yin89]. width [Lec85]. widths [Sch89b]. Wilcoxon [Bru88, TH87]. Wilk [SH87]. window [Sch89b]. Winsorization [VM83]. Wishart [Ric83]. within [KM87a, YR88].

yield [The87]. Yule [HI83, Kni89].

Zam [Fu89]. zero [LW82].

References

Anderson:1988:NCV

[AA88] Kevin K. Anderson and Krishna B. Athreya. A note on conjugate Π-variation and a weak limit theorem for the number of renewals. Statistics & Probability Letters, 6(3):151–154, February 1988. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715288901113.

Ahmed:1989:WAS

[AA89] A. N. Ahmed and A. A. Alzaid. Weak association with a stressstrength model application. Statistics & Probability Letters, 8(1): 91–95, May 1989. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715289900904.

Alzaid:1989:OPD

[AAO89] A. A. Alzaid and M. Al-Osh. Ordering probability distributions by tail behavior. Statistics & Probability Letters, 8(2):185– 188, June 1989. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715289900151.

Achcar:1986:LLM

[AB86] Jorge Alberto Achcar and Heleno Bolfarine. The log-linear model with a generalized gamma distribution for the error: a Bayesian approach. Statistics & Probability Letters, 4(6):325– 332, October 1986. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715286900532.

Arnold:1988:VBU

[AB88] Barry C. Arnold and Patrick L. Brockett. Variance bounds using a theorem of Pólya. Statistics & Probability Letters, 6(5): 321–326, April 1988. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715288900077.

Aihara:1989:IDP

[AB89] ShinIchi Aihara and Arunabha Bagchi. Infinite dimensional parameter identification for stochastic parabolic systems. Statistics & Probability Letters, 8(3):279–287, August 1989. CODEN SPLTDC.

ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016771528990134X.

Abdelhamid:1985:CRD

[Abd85] Sami N. Abdelhamid. On a characterization of rectangular distributions. Statistics & Probability Letters, 3(5):235–238, September 1985. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715285900379.

Abdous:1989:SUC

[Abd89] B. Abdous. Strong uniform consistency of kernel probability density estimators based on sample moments. Statistics & Probability Letters, 8(1):73–79, May 1989. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0167715289900874.

Abouammoh:1988:CMR

[Abo88] A. M. Abouammoh. On the criteria of the mean remaining life. Statistics & Probability Letters, 6(4):205–211, March 1988. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715288900624.

Anderson:1982:HFA

[AC82] John S. Anderson and Kenneth W. Clements. Holdings of financial assets: a Markov chain analysis. Statistics & Probability Letters, 1(1):36–40, July 1982. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0167715282900116.

Abouammoh:1986:CND

[AEN86] A. Abouammoh and E. El-Neweihi. Clusure of the NBUE and DMRL classes under formation of parallel systems. Statistics & Probability Letters, 4(5):223–225, August 1986. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715286900921.

Afsarinejad:1986:SOK

[Afs86] K. Afsarinejad. Self orthogonal Knut Vik designs. Statistics & Probability Letters, 4(6):289, October 1986. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715286900465.

Afsarinejad:1987:MOK

[Afs87] Kasra Afsarinejad. On mutually orthogonal Knut Vik designs. Statistics & Probability Letters, 5(5):323–324, August 1987. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715287900034.

Afsarinejad:1988:CBU

[Afs88a] Kasra Afsarinejad. Circular balanced uniform repeated measurements designs. Statistics & Probability Letters, 7(3):187–189, December 1988. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016771528890048X.

Afsarinejad:1988:SKV

[Afs88b] Kasra Afsarinejad. Semi Knut Vik designs. Statistics & Probability Letters, 6(4):243–245, March 1988. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715288900685.

Agresti:1983:SDP

[Agr83] Alan Agresti. A simple diagonals-parameter symmetry and quasisymmetry model. Statistics & Probability Letters, 1(6):313–316, October 1983. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715283900512.

Agresti:1989:AMK

[Agr89] Alan Agresti. An agreement model with kappa as parameter. Statistics & Probability Letters, 7(4):271–273, February 1989. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715289901041.

Arnold:1989:CPP

[AH89a] Barry C. Arnold and J. Terry Hallett. A characterization of the Pareto process among stationary stochastic processes of the form X_{n} = c min(X_{n−1}, Y_{n}). Statistics & Probability Letters, 8(4):377– 380, September 1989. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715289900473.

Aubuchon:1989:RBI

[AH89b] James C. Aubuchon and Thomas P. Hettmansperger. Rankbased inference for linear models: asymmetric errors. Statistics & Probability Letters, 8(2):97–107, June 1989. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715289900011.

Ahmad:1985:PAM

[Ahm85] Ibrahim A. Ahmad. On the Poisson approximation of multinomial probabilities. Statistics & Probability Letters, 3(1):55– 56, February 1985. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715285900136.

Antoch:1989:NRQ

[AJ89] J. Antoch and P. Janssen. Nonparametric regression M-quantiles. Statistics & Probability Letters, 8(4):355–362, September 1989. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0167715289900448.

Aras:1989:LDS

[AJZ89] Girish Aras, S. Rao Jammalamadaka, and X. Zhou. Limit distribution of spacings statistics when the sample size is random. Statistics & Probability Letters, 8(5):451–456, October 1989. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715289900266.

Azor:1983:SEM

[AK83] Meir Azor and Peter Kubat. Simple estimators for the mean life in destructive attribute life tests. Statistics & Probability Letters, 1(3):151–154, March 1983. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0167715283900640.

Ahmad:1988:TDO

[AK88] Ibrahim A. Ahmad and Subhash C. Kochar. Testing for dispersive ordering. Statistics & Probability Letters, 7(3):179–185, December 1988. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715288900478.

Akaike:1982:FLP

[Aka82] Hirotugu Akaike. On the fallacy of the likelihood principle. Statistics & Probability Letters, 1(2):75–78, November 1982. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0167715282900190.

Akritas:1982:NCN

[Akr82] Michael G. Akritas. A note on contiguity and L_{1}-norm. Statistics & Probability Letters, 1(1):23–25, July 1982. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715282900074.

Akritas:1984:SPC

[Akr84] Michael G. Akritas. A simple proof for the Chernoff–Savage theorem. Statistics & Probability Letters, 2(1):39–44, January 1984. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771528490035X.

Akritas:1988:ADN

[Akr88] Michael G. Akritas. An asymptotic derivation of Neyman’s C(α) test. Statistics & Probability Letters, 6(5):363–367, April 1988. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715288900144.

Albers:1984:EEU

[Alb84] Willem Albers. Efficient estimation under privacy restrictions in the disclosure problem. Statistics & Probability Letters, 2(1):1– 4, January 1984. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715284900270.

Albers:1988:RTR

[Alb88] Willem Albers. Rank tests for regression and k-sample rank tests under random censorship. Statistics & Probability Letters, 6(5): 315–319, April 1988. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715288900065.

Aldous:1989:SMT

[Ald89] David J. Aldous. Stein’s method in a two-dimensional coverage problem. Statistics & Probability Letters, 8(4):307–314, September 1989. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715289900370.

Averous:1989:TLD

[AM89] J. Averous and M. Meste. Tailweight and life distributions. Statistics & Probability Letters, 8(4):381–387, September 1989. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0167715289900485.

Anderson:1984:DBE

[And84] Robert F. Anderson. Dynamics of Bayes estimates for the rate of Poisson processes with gamma priors and convex loss. Statistics & Probability Letters, 2(3):147–157, May 1984. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715284900063.

Andrews:1986:SMB

[And86] Richard W. Andrews. A statistic for measuring the balance of a sample. Statistics & Probability Letters, 4(1):39–41, January 1986. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715286900374.

Anderson:1989:EEV

[And89] T. W. Anderson. Evaluation of the expected value of a determinant. Statistics & Probability Letters, 8(1):25–26, May 1989. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715289900771.

Anonymous:1982:EB

[Ano82a] Anonymous. Editorial board. Statistics & Probability Letters, 1 (1):iii, July 1982. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715282900013.

Anonymous:1982:TUP

[Ano82b] Anonymous. Toulouse 82 Université Paul Sabatier. Statistics & Probability Letters, 1(1):51, July 1982. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715282900141.

Anonymous:1983:AIV

[Ano83] Anonymous. Author index volume 1 (1983). Statistics & Probability Letters, 1(6):333–335, October 1983. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016771528390055X.

Anonymous:1984:AIV

[Ano84a] Anonymous. Author index volume 2 (1984). Statistics & Probability Letters, 2(6):373–375, December 1984. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715284900506.

Anonymous:1984:E

[Ano84b] Anonymous. Erratum. Statistics & Probability Letters, 2(4): 255, August 1984. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715284900257.

Anonymous:1985:AIV

[Ano85a] Anonymous. Author index volume 3 (1985). Statistics & Probability Letters, 3(6):343–345, October 1985. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715285900689.

Anonymous:1985:EB

[Ano85b] Anonymous. Editorial board. Statistics & Probability Letters, 3 (1):ii, February 1985. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016771528590001X.

Anonymous:1986:Eb

[Ano86a] Anonymous. Errata. Statistics & Probability Letters, 4(2): 107, March 1986. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016771528690026X.

Anonymous:1986:Ec

[Ano86b] Anonymous. Errata. Statistics & Probability Letters, 4(4): 217, June 1986. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715286900702.

Anonymous:1986:Ea

[Ano86c] Anonymous. Erratum. Statistics & Probability Letters, 4(1): 52, January 1986. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715286900416.

Anonymous:1986:SPC

[Ano86d] Anonymous. A simple proof for the Chernoff–Savage theorem by Michael G. Akritas. Statistics & Probability Letters, 4(6): 349, October 1986. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016771528690057X.

Anonymous:1987:AIV

[Ano87a] Anonymous. Author index volumes 4 and 5 (1986 and 1987). Statistics & Probability Letters, 5(6):429–433, October 1987. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0167715287900952.

Anonymous:1987:EB

[Ano87b] Anonymous. Editorial board. Statistics & Probability Letters, 6(1): ii, September 1987. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715287900484.

Anonymous:1987:Ec

[Ano87c] Anonymous. Errata. Statistics & Probability Letters, 5(2): 161, March 1987. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715287900733.

Anonymous:1987:Ea

[Ano87d] Anonymous. Erratum. Statistics & Probability Letters, 5(1): 81, January 1987. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715287900319.

Anonymous:1987:Ed

[Ano87e] Anonymous. Erratum. Statistics & Probability Letters, 5(4): 313, June 1987. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715287901118.

Anonymous:1987:Ee

[Ano87f] Anonymous. Erratum. Statistics & Probability Letters, 5(5): 379, August 1987. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715287900150.

Anonymous:1988:AIV

[Ano88a] Anonymous. Author index volume 6 (1988). Statistics & Probability Letters, 6(6):447–449, May 1988. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016771528890106X.

Anonymous:1988:EB

[Ano88b] Anonymous. Editorial board. Statistics & Probability Letters, 7(1): ii, July 1–87, 1988. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715288900764.

Anonymous:1988:Ea

[Ano88c] Anonymous. Errata. Statistics & Probability Letters, 6(5): 369, April 1988. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715288900156.

Anonymous:1988:Ec

[Ano88d] Anonymous. Errata. Statistics & Probability Letters, 7(1):87, July 1–87, 1988. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715288900934.

Anonymous:1988:Ed

[Ano88e] Anonymous. Erratum. Statistics & Probability Letters, 7(3):265, December 1988. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715288900612.

Anonymous:1989:AIVa

[Ano89a] Anonymous. Author index volume 7 (1988/1989). Statistics & Probability Letters, 7(5):441–443, April 1989. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715289901028.

Anonymous:1989:AIVb

[Ano89b] Anonymous. Author index volume 8 (1989). Statistics & Probability Letters, 8(5):497–499, October 1989. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715289900345.

Anonymous:1989:E

[Ano89c] Anonymous. Erratum. Statistics & Probability Letters, 7(4): 349, February 1989. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016771528990120X.

Anonymous:1989:ECT

[Ano89d] Anonymous. Erratum: “Convolution type estimators for nonparametric regression”, by Y. P. Mack and Hans-Georg Müller [Statistics & Probability Letters 7 (3) 229–239]. Statistics & Probability Letters, 8(2):195, June 1989. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0167715289900199. See [MM88a].

Athreya:1986:NSM

[AP86] Krishna B. Athreya and Sastry G. Pantula. A note on strong mixing of ARMA processes. Statistics & Probability Letters, 4(4): 187–190, June 1986. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715286900647.

Arenas:1989:PPP

[Are89] C. Arenas. On point processes in the plane. Statistics & Probability Letters, 8(4):325–328, September 1989. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715289900394.

Arnold:1985:SI

[Arn85] Steven F. Arnold. Sufficiency and invariance. Statistics & Probability Letters, 3(5):275–279, September 1985. CODEN SPLTDC.

ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715285900446.

Arnold:1987:BDP

[Arn87] Barry C. Arnold. Bivariate distributions with Pareto conditionals. Statistics & Probability Letters, 5(4):263–266, June 1987. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715287901027.

Arnold:1988:LPC

[Arn88] Barry C. Arnold. A logistic process constructed using geometric minimization. Statistics & Probability Letters, 7(3):253–257, December 1988. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715288900594.

Aguirre-Torres:1985:TSN

[AT85] Victor Aguirre-Torres. Testing several nonnested regressions simultaneously. a nonparametric approach. Statistics & Probability Letters, 3(6):317–324, October 1985. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715285900641.

Anantharam:1989:PMC

[AT89] V. Anantharam and P. Tsoucas. A proof of the Markov chain tree theorem. Statistics & Probability Letters, 8(2):189– 192, June 1989. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715289900163.

Athreya:1983:SLB

[Ath83] K. B. Athreya. Strong law for the bootstrap. Statistics & Probability Letters, 1(3):147–150, March 1983. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715283900639.

Athreya:1985:RRM

[Ath85] K. B. Athreya. On the range of recurrent Markov chains. Statistics & Probability Letters, 3(3):143–145, June 1985. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715285900525.

Athreya:1986:ACF

[Ath86] K. B. Athreya. Another conjugate family for the normal distribution. Statistics & Probability Letters, 4(2):61–64, March 1986. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715286900180.

Athreya:1987:MBF

[Ath87] Krishna B. Athreya. Modified Bessel function asymptotics via probability. Statistics & Probability Letters, 5(5):325–327, August 1987. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715287900046.

Arnold:1986:LOM

[AV86] Barry C. Arnold and Jose A. Villaseñor. Lorenz ordering of means and medians. Statistics & Probability Letters, 4(1):47– 49, January 1986. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715286900398.

Abaya:1984:SRE

[AW84] Efren F. Abaya and Gary L. Wise. Some remarks on the existence of optimal quantizers. Statistics & Probability Letters, 2 (6):349–351, December 1984. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0167715284900452.

Babu:1986:EER

[Bab86] Gutti Jogesh Babu. Efficient estimation of the reciprocal of the density quantile function at a point. Statistics & Probability Letters, 4(3):133–139, April 1986. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0167715286900763.

Bajorski:1987:LBO

[Baj87] Piotr Bajorski. Local Bahadur optimality of some rank tests of independence. Statistics & Probability Letters, 5(4):255– 262, June 1987. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715287901015.

Ballerini:1987:ACT

[Bal87] Rocco Ballerini. Another characterization of the type I extreme value distribution. Statistics & Probability Letters, 5(2): 83–85, March 1987. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715287900605.

Baltagi:1989:ANS

[Bal89] Badi H. Baltagi. Applications of a necessary and sufficient condition for OLS to be BLUE. Statistics & Probability Letters, 8 (5):457–461, October 1989. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0167715289900278.

Bartoszewicz:1986:DOT

[Bar86] Jaroslaw Bartoszewicz. Dispersive ordering and the total time on test transformation. Statistics & Probability Letters, 4(6):285– 288, October 1986. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715286900453.

Bartoszewicz:1987:NDO

[Bar87] J. Bartoszewicz. A note on dispersive ordering defined by hazard functions. Statistics & Probability Letters, 6(1):13–16, September 1987. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715287900526.

Berkane:1986:MED

[BB86] Maia Berkane and P. M. Bentler. Moments of elliptically distributed random variates. Statistics & Probability Letters, 4(6):333– 335, October 1986. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715286900544.

Berkane:1987:DKE

[BB87] Maria Berkane and P. M. Bentler. Distribution of kurtoses, with estimators and tests of homogeneity of kurtosis. Statistics & Probability Letters, 5(3):201–207, April 1987. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016771528790040X.

Babu:1988:BCI

[BB88] G. J. Babu and A. Bose. Bootstrap confidence intervals. Statistics & Probability Letters, 7(2):151–160, September 1988. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0167715288900429.

Birch:1989:CIR

[BB89] Nancy J. Birch and Richard K. Burdick. Confidence intervals on the ratio of expected mean squares (θ_{1} + θ_{2} + θ_{3})/θ_{4}. Statistics & Probability Letters, 7(4):335–342, February 1989. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715289901181.

Bonett:1989:ARC

[BBW89] Douglas G. Bonett, P. M. Bentler, and J. Arthur Woodward. Asymptotic results for a conditional Poisson log-linear model. Statistics & Probability Letters, 7(4):267–270, February 1989. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
016771528990103X.

Basu:1983:RSS

[BE83] A. P. Basu and Nader Ebrahimi. On the reliability of stochastic systems. Statistics & Probability Letters, 1(5):265–267, August 1983. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715283900421.

Basu:1986:MNB

[BE86] A. P. Basu and Nader Ebrahimi. Multivariate new better than used in expectation distributions. Statistics & Probability Letters, 4(6):295–301, October 1986. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0167715286900489.

Becker:1989:BND

[Bec89] Mark P. Becker. On the bivariate normal distribution and association models for ordinal categorical data. Statistics & Probability Letters, 8(5):435–440, October 1989. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715289900230.

Berger:1983:CSP

[Ber83] Agnes Berger. On comparing survival probabilities from discrete observations under unequal censoring. Statistics & Probability Letters, 1(5):251–257, August 1983. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0167715283900408.

Berman:1989:ASB

[Ber89] Mark Berman. The asymptotic statistical behaviour of some estimators for a circular structural model. Statistics & Probability Letters, 7(5):413–416, April 1989. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0167715289900965.

Bock:1988:NNC

[BG88a] M. E. Bock and Z. Govindarajulu. A note on the noncentral chi-square distribution. Statistics & Probability Letters, 7(2):127– 129, September 1988. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715288900375.

Burke:1988:GFB

[BG88b] Murray D. Burke and Edit Gombay. On goodness-of-fit and the bootstrap. Statistics & Probability Letters, 6(5):287–293, April 1988. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715288900016.

Booth:1986:GSE

[BGM^{+}86] J. Booth, J. Gani, M.-P. Malice, H. Mansouri, and G. Maravankin. A general solution for the epidemic with carriers. Statistics & Probability Letters, 4(1):9–15, January 1986. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715286900313.

Baksalary:1984:IIC

[BH84] Jerzy K. Baksalary and Jan Hauke. Inheriting independence and chi-squaredness under certain matrix orderings. Statistics & Probability Letters, 2(1):35–38, January 1984. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715284900348.

Bohning:1986:MCS

[BH86a] D. Böhning and H. Holling. On minimizing chi-square distances under the hypothesis of homogeneity or independence for a twoway contingency table. Statistics & Probability Letters, 4(5):253– 258, August 1986. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715286900982.

Burke:1986:EIF

[BH86b] Murray D. Burke and Lajos Horváth. Estimation of influence functions. Statistics & Probability Letters, 4(2):81–85, March 1986. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715286900222.

Bhave:1987:GPA

[Bha87] S. V. Bhave. Godambe’s paradox and the ancillarity principle. Statistics & Probability Letters, 5(4):243–246, June 1987. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0167715287900988.

Bhave:1989:IAL

[Bha89] S. V. Bhave. Inconsistencies in the argument leading to the rule of succession. Statistics & Probability Letters, 7(5):435– 440, April 1989. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715289901016.

Birkel:1988:NSL

[Bir88] Thomas Birkel. A note on the Strong Law of Large Numbers for positively dependent random variables. Statistics & Probability Letters, 7(1):17–20, July 1–87, 1988. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0167715288900806.

Bar-Lev:1987:EMA

[BLE87] Shaul K. Bar-Lev and Peter Enis. Existence of moments and an asymptotic result based on a mixture of exponential distributions. Statistics & Probability Letters, 5(4):273–277, June 1987. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715287901040.

Bera:1987:EAT

[BM87] Anil K. Bera and Michael McAleer. On exact and asymptotic tests of non-nested models. Statistics & Probability Letters, 5(1): 19–22, January 1987. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715287900204.

Boker:1988:SRM

[Bök88] Fred Böker. Some remarks on measures of association of 2 × 2 tables. Statistics & Probability Letters, 6(6):395–397, May 1988. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715288900983.

Bolfarine:1989:PVP

[Bol89] Heleno Bolfarine. Population variance prediction under normal dynamic superpopulation models. Statistics & Probability Letters, 8 (1):35–39, May 1989. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715289900801.

Bonett:1985:LNM

[Bon85] Douglas G. Bonett. A linear negative multinomial model. Statistics & Probability Letters, 3(3):127–129, June 1985. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715285900495.

Bonett:1986:MEC

[Bon86] Douglas G. Bonett. Mixed estimation in covariance structure models. Statistics & Probability Letters, 4(4):157–159, June 1986. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715286900581.

Bonett:1989:PCS

[Bon89] Douglas G. Bonett. Pearson chi-square estimator and test for log-linear models with expected frequencies subject to linear constraints. Statistics & Probability Letters, 8(2):175–177, June 1989. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715289900138.

Berger:1985:SIC

[BP85] Agnes Berger and Ora E. Percus. On sampling by index cases. Statistics & Probability Letters, 3(4):213–219, July 1985. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0167715285900215.

Boland:1987:SCM

[BP87] Philip J. Boland and Frank Proschan. Schur convexity of the maximum likelihood function for the multivariate hypergeometric and multinomial distributions. Statistics & Probability Letters, 5 (5):317–322, August 1987. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0167715287900022.

Best:1988:TBN

[BR88] D. J. Best and J. C. W. Rayner. A test for bivariate normality. Statistics & Probability Letters, 6(6):407–412, May 1988. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715288901009.

Bradley:1983:ANS

[Bra83] Richard C. Bradley. Asymptotic normality of some kernel-type estimators of probability density. Statistics & Probability Letters, 1 (6):295–300, October 1983. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0167715283900482.

Bradley:1989:CMC

[Bra89] Richard C. Bradley. A caution on mixing conditions for random fields. Statistics & Probability Letters, 8(5):489–491, October 1989. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715289900321.

Bross:1982:SCC

[Bro82] Irwin D. J. Bross. Simplicity and credibility: a counterstrategy. Statistics & Probability Letters, 1(2):79–83, November 1982. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715282900207.

Brus:1988:RFD

[Bru88] Tomek Brus. A recurrence formula for the distribution of the Wilcoxon rank sum statistic. Statistics & Probability Letters, 7 (2):161–165, September 1988. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0167715288900430.

Barigelli:1984:RRC

[BS84] Bruno Barigelli and Romana Scozzafava. Remarks on the role of conditional probability in data exploration. Statistics & Probability Letters, 2(1):15–18, January 1984. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0167715284900300.

Block:1985:CND

[BSS85] Henry W. Block, Thomas H. Savits, and Moshe Shaked. A concept of negative dependence using stochastic ordering. Statistics & Probability Letters, 3(2):81–86, April 1985. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016771528590029X.

Bertin:1984:SCD

[BT84] Emile M. J. Bertin and Radu Theodorescu. Some characterizations of discrete unimodality. Statistics & Probability Letters, 2 (1):23–30, January 1984. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0167715284900324.

Bueno:1988:ICM

[Bue88] Vanderlei Costa Bueno. On the importance of components for multistate monotone systems. Statistics & Probability Letters, 7 (1):51–59, July 1–87, 1988. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0167715288900880.

Burton:1985:PPC

[Bur85] Robert M. Burton. Pointwise properties of convergence in probability. Statistics & Probability Letters, 3(6):315–316, October 1985. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771528590063X.

Burman:1989:RCS

[Bur89] Prabir Burman. Rate of convergence of the spline estimates for Markov chains. Statistics & Probability Letters, 8(3):245– 253, August 1989. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715289901296.

Berger:1989:TT

[BW89] Agnes Berger and Sylvan Wallenstein. On the theory of C_{α}-tests. Statistics & Probability Letters, 7(5):419–424, April 1989. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0167715289900989.

Cacoullos:1983:CMD

[Cac83] T. Cacoullos. Characterizations of mixtures of discrete distributions by a regression point. Statistics & Probability Letters, 1 (5):269–272, August 1983. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0167715283900433.

Cacoullos:1987:CGM

[Cac87] T. Cacoullos. Characterizations of generalized multivariate discrete distributions by a regression point. Statistics & Probability Letters, 5(1):39–42, January 1987. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0167715287900241.

Chaudhuri:1989:TWP

[CC89] Arijit Chaudhuri and Nimai Kumar Chandra. A test for Weibull populations. Statistics & Probability Letters, 7(5):377–380, April 1989. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715289900552.

Chernick:1982:EAP

[CD82] Michael R. Chernick and Richard A. Davis. Extremes in autoregressive processes with uniform marginal distributions. Statistics & Probability Letters, 1(2):85–88, November 1982. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715282900219.

Chakraborti:1988:CDF

[CD88] S. Chakraborti and M. M. Desu. A class of distribution-free tests for testing homogeneity against ordered alternatives. Statistics & Probability Letters, 6(4):251–256, March 1988. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715288900703.

Cheng:1983:AOC

[CF83] Smiley W. Cheng and James C. Fu. An algorithm to obtain the critical values of the t, χ^{2} and F distributions. Statistics & Probability Letters, 1(5):223–227, August 1983. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715283900354.

Cheng:1986:FOM

[CF86] Ping Cheng and James C. Fu. On a fundamental optimality of the maximum likelihood estimator. Statistics & Probability Letters, 4(4):173–178, June 1986. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0167715286900611.

Cressie:1984:MBC

[CG84] Noel Cressie and Gary Glonek. Median based covariogram estimators reduce bias. Statistics & Probability Letters, 2(5):299– 304, October 1984. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715284900695.

Costantini:1983:CSS

[CGN83] C. Costantini, A. Gerardi, and G. Nappo. On the convergence of sequences of stationary jump Markov processes. Statistics & Probability Letters, 1(3):155–160, March 1983. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715283900652.

Csorgo:1986:AWE

[CH86] Miklós Csörgő and Lajos Horváth. Approximations of weighted empirical and quantile processes. Statistics & Probability Letters, 4 (6):275–280, October 1986. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.sciencedirect.

com/science/article/pii/016771528690043X.

Carroll:1989:SNN

[CH89a] R. J. Carroll and W. Härdle. Symmetrized nearest neighbor regression estimates. Statistics & Probability Letters, 7(4):315– 318, February 1989. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715289901144.

Cheng:1989:PBE

[CH89b] Kan Cheng and Zongfu He. On proximity between exponential and DMRL distributions. Statistics & Probability Letters, 8(1): 55–57, May 1989. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715289900837.

Csorgo:1989:BPA

[CH89c] Miklós Csörgő and Lajos Horváth. On best possible approximations of local time. Statistics & Probability Letters, 8(4):301–306, September 1989. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715289900369.

Chanda:1985:SDC

[Cha85a] Kamal C. Chanda. Sampling distribution for a class of estimators for nonregular linear processes. Statistics & Probability Letters, 3 (5):261–268, September 1985. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0167715285900422.

Chaubey:1985:ECT

[Cha85b] Yogendra P. Chaubey. Estimators of covariances in time series models. Statistics & Probability Letters, 3(1):51–53, February 1985. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715285900124.

Chao:1986:EBP

[Cha86] Min-Te Chao. On M and P estimators that have breakdown point equal to 12. Statistics & Probability Letters, 4(3):127– 131, April 1986. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715286900751.

Chawla:1988:ETG

[Cha88] J. S. Chawla. The existence theorem in general ridge regression. Statistics & Probability Letters, 7(2):135–137, September 1988. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715288900399.

Chassaing:1989:ORN

[Cha89] Philippe Chassaing. An optimal random number generator on Z_{p}. Statistics & Probability Letters, 7(4):307–309, February 1989. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715289901120.

Cheng:1984:HTL

[Che84] Smiley W. Cheng. Hypothesis testing of the location and scale parameters using order statistics. Statistics & Probability Letters, 2 (4):207–210, August 1984. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0167715284900166.

Chiu:1989:BSK

[Chi89] Shean-Tsong Chiu. Bandwidth selection for kernel estimate with correlated noise. Statistics & Probability Letters, 8(4):347–354, September 1989. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715289900436.

Csorgo:1986:HLM

[CHR86] Miklós Csörgő, Lajos Horváth, and Pál Révész. How large must be the difference between local time and mesure du voisinage of Brownian motion? Statistics & Probability Letters, 4(4):161– 166, June 1986. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715286900593.

Chib:1987:NDP

[CJT87] Siddhartha Chib, S. Rao Jammalamadaka, and Ram C. Tiwari. A new definition of the predictive likelihood. Statistics & Probability Letters, 5(2):113–118, March 1987. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715287900654.

Ceranka:1987:RBB

[CK87] B. Ceranka and K. Katulska. A relation between BIB designs and chemical balance weighing designs. Statistics & Probability Letters, 5(5):339–341, August 1987. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0167715287900071.

Clarotti:1988:FMB

[CKS88] C. A. Clarotti, G. Koch, and F. Spizzichino. A filtering model for Bayesian analysis of failure data contaminated by maintenance. Statistics & Probability Letters, 7(2):145–150, September 1988. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715288900417.

Chhetry:1986:DSW

[CKZ86] Devendra Chhetry, George Kimeldorf, and Hassan Zahedi. Dependence structures in which uncorrelatedness implies independence. Statistics & Probability Letters, 4(4):197–201, June 1986. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715286900660.

Cheng:1987:MLE

[CL87] Philip E. Cheng and Gwo Dong Lin. Maximum likelihood estimation of a survival function under the koziol-green proportional hazards model. Statistics & Probability Letters, 5(1):75– 80, January 1987. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715287900307.

Clarke:1989:UMD

[Cla89] Brenton R. Clarke. An unbiased minimum distance estimator of the proportion parameter in a mixture of two normal distributions. Statistics & Probability Letters, 7(4):275–281, February 1989. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715289901053.

Chou:1984:DSS

[CM84] Youn-Min Chou and Robert W. Mee. Determination of sample sizes for setting β-content tolerance limits controlling both tails of the normal distribution. Statistics & Probability Letters, 2(5):311– 314, October 1984. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715284900713.

Csorgo:1988:DES

[CM88a] Sándor Csörgö and Jan Mielniczuk. Density estimation in the simple proportional hazards model. Statistics & Probability Letters, 6(6):419–426, May 1988. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0167715288901022.

Cuesta:1988:UCM

[CM88b] Juan A. Cuesta and Carlos Matrán. Uniform consistency of rmeans. Statistics & Probability Letters, 7(1):65–71, July 1–87, 1988. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715288900909.

Cooil:1988:WIP

[Coo88] Bruce Cooil. When are intermediate processes of the same stochastic order? Statistics & Probability Letters, 6(3):159–162, February 1988. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715288901137.

Cox:1983:NNN

[Cox83] Trevor F. Cox. Nearest neighbours to nearest neighbours. Statistics & Probability Letters, 1(4):161–166, June 1983. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715283900238.

Cacoullos:1985:UBV

[CP85] T. Cacoullos and V. Papathanasiou. On upper bounds for the variance of functions of random variables. Statistics & Probability Letters, 3(4):175–184, July 1985. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0167715285900148.

Cacoullos:1986:BVF

[CP86] T. Cacoullos and V. Papathanasiou. Bounds for the variance of functions of random variables by orthogonal polynomials and Bhattacharya bounds. Statistics & Probability Letters, 4(1):21– 23, January 1986. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715286900337.

Cacoullos:1989:CDV

[CP89] T. Cacoullos and V. Papathanasiou. Characterizations of distributions by variance bounds. Statistics & Probability Letters, 7(5): 351–356, April 1989. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715289900503.

Chen:1986:BDB

[CR86] Jeesen Chen and Herman Rubin. Bounds for the difference between median and mean of gamma and Poisson distributions. Statistics & Probability Letters, 4(6):281–283, October 1986. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715286900441.

Chanda:1989:AEP

[CR89] Kamal C. Chanda and F. H. Ruymgaart. Asymptotic estimate of probability of misclassification for discriminant rules based on density estimates. Statistics & Probability Letters, 8(1):81– 88, May 1989. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715289900886.

Carroll:1985:NLT

[CS85] Raymond J. Carroll and Helmut Schneider. A note on Levene’s tests for equality of variances. Statistics & Probability Letters, 3(4):191–194, July 1985. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0167715285900161.

Choi:1987:IS

[CS87a] Bong Dae Choi and Soo Hak Sung. On the integrability of sup |S_{n}|/n^{1}/r for 1 < r < 2. Statistics & Probability Letters, 5(4):267–272, June 1987. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0167715287901039.

Cohen:1987:AGF

[CS87b] Arthur Cohen and Harold B. Sackrowitz. Admissibility of goodness of fit tests for discrete exponential families. Statistics & Probability Letters, 5(1):1–3, January 1987. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0167715287900174.

Chow:1988:NPE

[CS88] Shein-Chung Chow and Jun Shao. A new procedure for the estimation of variance components. Statistics & Probability Letters, 6(5):349–355, April 1988. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0167715288900120.

Cohen:1989:TSC

[CS89] Arthur Cohen and Harold B. Sackrowitz. Two stage conditionally unbiased estimators of the selected mean. Statistics & Probability Letters, 8(3):273–278, August 1989. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715289901338.

Csorgo:1985:TL

[Csö85] Sándor Csörgő. Testing for linearity. Statistics & Probability Letters, 3(1):45–49, February 1985. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0167715285900112.

Conway:1982:CME

[CT82] Delores Conway and Henri Theil. The conditioning of the maximum entropy covariance matrix and its inverse. Statistics & Probability Letters, 1(2):103–106, November 1982. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715282900244.

Caselton:1984:OMN

[CZ84] W. F. Caselton and J. V. Zidek. Optimal monitoring network designs. Statistics & Probability Letters, 2(4):223–227, August 1984. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715284900208.

Dabrowski:1982:NTB

[Dab82] AndréRobert Dabrowski. A note on a theorem of Berkes and Philipp for dependent sequences. Statistics & Probability Letters, 1(2):53–55, November 1982. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0167715282900153.

Dabrowski:1985:FLI

[Dab85] AndréRobert Dabrowski. A functional law of the iterated logarithm for associated sequences. Statistics & Probability Letters, 3(4):209–212, July 1985. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0167715285900203.

Dabrowski:1988:STI

[Dab88] AndréRobert Dabrowski. Strassen-type invariance principles for exchangeable sequences. Statistics & Probability Letters, 7(1):23– 26, July 1–87, 1988. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016771528890082X.

Dabrowski:1989:PIL

[Dab89] Jacek Dabrowski. Parameter identification in linear stochastic differential equations. Statistics & Probability Letters, 7(5):391– 394, April 1989. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715289900928.

Dalen:1987:ABS

[Dal87] Jörgen Dalén. Algebraic bounds on standardized sample moments. Statistics & Probability Letters, 5(5):329–331, August 1987. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715287900058.

Dale:1988:EOP

[Dal88] A. I. Dale. An early occurrence of the Poisson distribution. Statistics & Probability Letters, 7(1):21–22, July 1–87, 1988. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715288900818.

Danial:1988:GSC

[Dan88] Edward J. Danial. Generalization to the sufficient conditions for a discrete random variable to be infinitely divisible. Statistics & Probability Letters, 6(6):379–382, May 1988. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715288900958.

Danial:1989:CSD

[Dan89] Edward J. Danial. Characterization of the special discrete distributions through infinite, noninfinite and finite divisibility. Statistics & Probability Letters, 8(1):1–7, May 1989. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715289900734.

Davis:1986:RBS

[Dav86] Linda June Davis. Relationship between strictly collapsible and perfect contingency tables. Statistics & Probability Letters, 4(3): 119–122, April 1986. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715286900738.

Davis:1987:PCM

[Dav87] Linda June Davis. Partial collapsibility in multidimensional contingency tables. Statistics & Probability Letters, 5(2):129– 134, March 1987. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016771528790068X.

Deheuvels:1989:NPE

[DdO89] Paul Deheuvels and José Tiago de Oliveira. On the nonparametric estimation of the bivariate extreme-value distributions. Statistics & Probability Letters, 8(4):315–323, September 1989. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715289900382.

Dehling:1989:CCT

[Deh89] Herold Dehling. Complete convergence of triangular arrays and the law of the iterated logarithm for U-statistics. Statistics & Probability Letters, 7(4):319–321, February 1989. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715289901156.

Devroye:1984:MGR

[Dev84] Luc Devroye. Methods for generating random variates with Pólya characteristic functions. Statistics & Probability Letters, 2(5):257– 261, October 1984. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715284900610.

Devroye:1989:NCC

[Dev89a] Luc Devroye. On the non-consistency of the L_{2}-cross-validated kernel density estimate. Statistics & Probability Letters, 8(5):425– 433, October 1989. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715289900229.

Devroye:1989:ULB

[Dev89b] Luc Devroye. A universal lower bound for the kernel estimate. Statistics & Probability Letters, 8(5):419–423, October 1989. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715289900217.

Dey:1987:IEM

[Dey87] Dipak K. Dey. Improved estimation of a multinormal precision matrix. Statistics & Probability Letters, 6(2):125–128, November 1987. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016771528790085X.

Diaz:1988:VPA

[DG88] M. Martin Diaz and B. Salvador González. The validity of the “Pool-Adjacent-Violator” algorithm. Statistics & Probability Letters, 6 (3):143–145, February 1988. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0167715288901095.

Dorsett:1983:MEW

[DGG83] Dovalee Dorsett, Richard F. Gunst, and Eugene C. Gartland, Jr. Multicollinear effects of weighted least squares regression. Statistics & Probability Letters, 1(4):207–211, June 1983. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715283900329.

Dia:1987:ERF

[Dia87] Galaye Dia. Estimation of a regression function on a Poisson process. Statistics & Probability Letters, 6(1):47–54, September 1987. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715287900587.

Denker:1989:SLD

[DJ89] Manfred Denker and Adam Jakubowski. Stable limit distributions for strongly mixing sequences. Statistics & Probability Letters, 8 (5):477–483, October 1989. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0167715289900308.

Devroye:1986:LBI

[DLS86] Luc Devroye, Gerard Letac, and Vanamamalai Seshadri. The limit behavior of an interval splitting scheme. Statistics & Probability Letters, 4(4):183–186, June 1986. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0167715286900635.

DeDominicis:1984:CPA

[DM84] R. De Dominicis and R. Manca. A computational procedure for the asymptotic analysis of homogeneous semi-Markov processes. Statistics & Probability Letters, 2(4):249–253, August 1984. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715284900245.

deOliveira:1989:IED

[dO89] J. Tiago de Oliveira. Intrinsic estimation of the dependence structure for bivariate extremes. Statistics & Probability Letters, 8 (3):213–218, August 1989. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0167715289901247.

delPino:1984:LRT

[dP84] Guido E. del Pino. Linear restrictions and two step least squares with applications. Statistics & Probability Letters, 2(4):245– 248, August 1984. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715284900233.

delPino:1985:RLE

[dP85] Guido E. del Pino. On restricted linear estimation for regression in stochastic processes. Statistics & Probability Letters, 3(1):9– 13, February 1985. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715285900033.

delPino:1987:SOP

[dP87] Guido E. del Pino. Second order processes with restrictions in the index set and applications. Statistics & Probability Letters, 6(1):1–5, September 1987. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0167715287900496.

Dandawate:1984:CKS

[DPV84] P. N. Dandawate, G. A. Patwardhan, and M. N. Vartak. Connectedness of Kronecker sum and partial Kronecker row sum of designs. Statistics & Probability Letters, 2(6):315–317, December 1984. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715284900385.

delRio:1988:CSD

[dR88] Manuel del Río. Cook statistic and diagnostic for transformations. Statistics & Probability Letters, 6(4):281–286, March 1988. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715288900752.

Dehling:1988:FLI

[DT88] Herold Dehling and Murad S. Taqqu. The functional law of the iterated logarithm for the empirical process of some long-range dependent sequences. Statistics & Probability Letters, 7(1):81–85, July 1–87, 1988. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715288900922.

Eberlein:1984:WCP

[Ebe84] Ernst Eberlein. Weak convergence of partial sums of absolutely regular sequences. Statistics & Probability Letters, 2(5):291–293, October 1984. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715284900671.

Ebrahimi:1985:TSH

[Ebr85] Nader Ebrahimi. Two suggestions of how to define a stochastic stress-strength system. Statistics & Probability Letters, 3(6):295– 297, October 1985. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715285900586.

Edler:1984:FBP

[Edl84] Lutz Edler. Fluctuation behavior of Poisson triangular arrays. Statistics & Probability Letters, 2(4):203–205, August 1984. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715284900154.

Eubank:1986:DPL

[EG86] R. L. Eubank and R. F. Gunst. Diagnostics for penalized leastsquares estimators. Statistics & Probability Letters, 4(5):265– 272, August 1986. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016771528690101X.

Einmahl:1987:GFL

[Ein87] John H. J. Einmahl. A general form of the law of the iterated logarithm for the weighted multivariate empirical process. Statistics & Probability Letters, 5(3):181–185, April 1987. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715287900356.

Elliott:1988:SPM

[EK88] Robert J. Elliott and Michael Kohlmann. A short proof of a martingale representation result. Statistics & Probability Letters, 6(5):327–329, April 1988. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0167715288900089.

El-Neweihi:1984:CCU

[EN84] Emad El-Neweihi. Characterizations and closure under convolution of two classes of multivariate life distributions. Statistics & Probability Letters, 2(6):333–335, December 1984. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715284900427.

Einmahl:1987:ASB

[ER87] J. H. J. Einmahl and F. H. Ruymgaart. The almost sure behavior of the oscillation modulus of the multivariate empirical process. Statistics & Probability Letters, 6(2):87–96, November 1987. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715287900794.

Eubank:1982:QDP

[Eub82] R. L. Eubank. A quantile domain perspective on the relationships between optimal grouping, spacing and stratification problems. Statistics & Probability Letters, 1(2):69–73, November 1982. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715282900189.

Eubank:1984:MSS

[Eub84] R. L. Eubank. The hat matrix for smoothing splines. Statistics & Probability Letters, 2(1):9–14, January 1984. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715284900294.

Evans:1989:ECL

[Eva89] Michael Evans. An example concerning the likelihood function. Statistics & Probability Letters, 7(5):417–418, April 1989. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0167715289900977.

Everitt:1988:FMM

[Eve88] B. S. Everitt. A finite mixture model for the clustering of mixed-mode data. Statistics & Probability Letters, 6(5):305– 309, April 1988. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715288900041.

Fabian:1988:LAM

[Fab88] Václav Fabian. The local asymptotic minimax adaptive property of a recursive estimate. Statistics & Probability Letters, 6(6): 383–388, May 1988. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016771528890096X.

Falk:1986:EQD

[Fal86] Michael Falk. On the estimation of the quantile density function. Statistics & Probability Letters, 4(2):69–73, March 1986. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0167715286900209.

Famoye:1987:SNG

[Fam87] Felix Famoye. A short note on the generalized logarithmic series distribution. Statistics & Probability Letters, 5(5):315–316, August 1987. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715287900010.

Fries:1986:ODI

[FB86] Arthur Fries and Gouri K. Bhattacharyya. Optimal design for an inverse Gaussian regression model. Statistics & Probability Letters, 4(6):291–294, October 1986. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0167715286900477.

Ferreiro:1987:MEM

[Fer87] Osvaldo Ferreiro. Methodologies for the estimation of missing observations in time series. Statistics & Probability Letters, 5(1):65– 69, January 1987. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715287900289.

Feuerverger:1988:BEN

[Feu88] Andrey Feuerverger. A bound for estimation in nonlinear time series models by independence testing methods. Statistics & Probability Letters, 6(4):237–241, March 1988. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715288900673.

Fang:1989:CMN

[FF89] Bi-Qi Fang and Kai-Tai Fang. A characterization of multivariate l_{1}-norm symmetric distributions. Statistics & Probability Letters, 7 (4):297–299, February 1989. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0167715289901090.

Finke:1984:MIE

[FFT84] Renate Finke, Lennart R. Flood, and Henri Theil. Minimum information estimation of allocation models of different sizes. Statistics & Probability Letters, 2(5):279–283, October 1984. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715284900658.

Feuerverger:1982:EMP

[FGS82] A. Feuerverger, I. Guttman, and S. K. Sinha. Estimation of a mean in the presence of an extreme observation. Statistics & Probability Letters, 1(2):89–96, November 1982. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715282900220.

Finster:1987:AFS

[Fin87] Mark P. Finster. An analysis of five simulation methods for determining the number of replications in a complex Monte Carlo study. Statistics & Probability Letters, 5(5):353–360, August 1987. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715287900101.

Fu:1984:NIB

[FK84] James C. Fu and Robert E. Kass. A note on the interpretation of the Bahadur bound and the rate of convergence of the maximum likelihood estimator. Statistics & Probability Letters, 2 (5):269–273, October 1984. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0167715284900634.

Fiebig:1983:SEE

[FKT83] Denzil G. Fiebig, Sartaj A. Kidwai, and Henri Theil. Simultaneous equation estimation from undersized samples. Statistics & Probability Letters, 1(5):229–232, August 1983. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715283900366.

Flury:1986:PCM

[Flu86a] Bernhard K. Flury. Proportionality of k covariance matrices. Statistics & Probability Letters, 4(1):29–33, January 1986. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0167715286900350.

Flury:1986:ATR

[Flu86b] Bernhard N. Flury. An asymptotic test for redundancy of variables in the comparison of two covariance matrices. Statistics & Probability Letters, 4(3):123–126, April 1986. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016771528690074X.

Fortney:1987:BAR

[FM87] William G. Fortney and Robert B. Miller. Bayesian analysis in random coefficient m-group regression. Statistics & Probability Letters, 5(2):135–142, March 1987. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0167715287900691.

Franklin:1988:ETS

[Fra88] LeRoy A. Franklin. Exact tables of Spearman’s Footrule for N = 11(1)18 with estimate of convergence and errors for the normal approximation. Statistics & Probability Letters, 6(6):399– 406, May 1988. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715288900995.

Flood:1984:AEE

[FRT84] Lennart R. Flood, Mercedes C. Rosalsky, and Henri Theil. Aggregation and the estimation of equation systems. Statistics & Probability Letters, 2(4):187–191, August 1984. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715284900129.

Feuerverger:1985:CQS

[FS85] Andrey Feuerverger and J. Michael Steele. On a characterization question for symmetric random variables. Statistics & Probability Letters, 3(1):35–37, February 1985. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715285900094.

Fu:1983:UBS

[Fu83] James C. Fu. On an upper bound of Sanov’s inequality for the probability of a large deviation. Statistics & Probability Letters, 1(3):107–113, March 1983. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0167715283900561.

Fu:1985:ERL

[Fu85] James C. Fu. On exponential rates of likelihood ratio estimators for location parameters. Statistics & Probability Letters, 3 (2):101–105, April 1985. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0167715285900343.

Fu:1989:MKZ

[Fu89] James C. Fu. Method of Kim–Zam: an algorithm for computing the maximum likelihood estimator. Statistics & Probability Letters, 8(3):289–296, August 1989. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0167715289901351.

Gajek:1987:IDF

[Gaj87] Lesław Gajek. On improving distribution function estimators which are not monotonic functions. Statistics & Probability Letters, 5(3): 213–215, April 1987. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715287900423.

Gather:1988:CED

[Gat88] Ursula Gather. On a characterization of the exponential distribution by properties of order statistics. Statistics & Probability Letters, 7 (2):93–96, September 1988. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0167715288900326.

Green:1986:REL

[GC86] Stephanie J. Green and John J. Crowley. On robust estimation of location for arbitrarily right-censored data. Statistics & Probability Letters, 4(6):303–308, October 1986. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715286900490.

Georgiev:1984:KEF

[Geo84] Alexander A. Georgiev. Kernel estimates of functions and their derivatives with applications. Statistics & Probability Letters, 2 (1):45–50, January 1984. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0167715284900361.

Georgiev:1988:APM

[Geo88] Alexander A. Georgiev. Asymptotic properties of the multivariate Nadaraya–Watson regression function estimate: The fixed design case. Statistics & Probability Letters, 7(1):35–40, July 1–87, 1988. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715288900855.

Gil:1989:SIM

[GG89] María Angeles Gil and Pedro Gil. On some information measures of degree β = 2: estimation in simple-stage cluster sampling. Statistics & Probability Letters, 8(2):157–162, June 1989. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715289900096.

Gut:1984:WTG

[GH84] Allan Gut and Lars Holst. On the waiting time in a generalized roulette game. Statistics & Probability Letters, 2(4):229– 239, August 1984. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016771528490021X.

Ghosh:1988:TMI

[Gho88] Subir Ghosh. On two methods of identifying influential sets of observations. Statistics & Probability Letters, 7(3):241–245, December 1988. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715288900570.

Gastwirth:1988:THS

[GJ88a] Joseph L. Gastwirth and Wesley O. Johnson. Testing the homogeneity of small error rates: application to the sensitivity of different Elisa tests for aids antibodies. Statistics & Probability Letters, 7(3):225–228, December 1988. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0167715288900557.

Guerrero:1988:TGB

[GJ88b] Victor M. Guerrero and Richard A. Johnson. Transforming grouped bivariate data to near normality. Statistics & Probability Letters, 6(4):213–224, March 1988. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0167715288900636.

Galambos:1989:SPC

[GK89] Janos Galambos and Imre Kátai. A simple proof for the continuity of infinite convolutions of binary random variables. Statistics & Probability Letters, 7(5):369–370, April 1989. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715289900539.

Gardiner:1989:PMI

[GL89] Joseph C. Gardiner and Shlomo Levental. On pricing of marketindexed certificates of deposit. Statistics & Probability Letters, 8 (4):329–334, September 1989. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0167715289900400.

Gupta:1985:RBP

[GM85] Ramesh C. Gupta and Joel E. Michalek. On the relationship between the proportional hazards and accelerated failure time models in survival analysis. Statistics & Probability Letters, 3(5):231– 234, September 1985. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715285900367.

Gatsonis:1987:ERN

[GMS87] Constantine Gatsonis, Brenda MacGibbon, and William Strawderman. On the estimation of a restricted normal mean. Statistics & Probability Letters, 6(1):21–30, September 1987. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016771528790054X.

Godambe:1987:RGP

[God87a] V. P. Godambe. Resolution of Godambe’s Paradox. Statistics & Probability Letters, 5(4):239, June 1987. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715287900964.

Godbole:1987:KSC

[God87b] Anant P. Godbole. On Klass’ series criterion for the minimal growth rate of partial maxima. Statistics & Probability Letters, 5(3):235–237, April 1987. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0167715287900472.

Goldstein:1988:CSS

[Gol88] Larry Goldstein. On the choice of step size in the Robbins– Monro procedure. Statistics & Probability Letters, 6(5):299– 303, April 1988. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016771528890003X.

Gouet:1989:EEP

[Gou89a] Raúl Gouet. Embedding in extremal processes and the asymptotic behavior of sums of minima. Statistics & Probability Letters, 8(3):219–223, August 1989. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0167715289901259.

Gouet:1989:MAS

[Gou89b] Raúl Gouet. A martingale approach to strong convergence in a generalized Pólya–Eggenberger urn model. Statistics & Probability Letters, 8(3):225–228, August 1989. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0167715289901260.

Greblicki:1985:PCH

[GP85] Włodzimierz Greblicki and Miroslaw Pawlak. Pointwise consistency of the Hermite series density estimate. Statistics & Probability Letters, 3(2):65–69, April 1985. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0167715285900264.

Genest:1989:CGF

[GR89] Christian Genest and Louis-Paul Rivest. A characterization of Gumbel’s family of extreme value distributions. Statistics & Probability Letters, 8(3):207–211, August 1989. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715289901235.

Grill:1988:ARW

[Gri88] Karl Grill. On the average of a random walk. Statistics & Probability Letters, 6(5):357–361, April 1988. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715288900132.

George:1987:ADB

[GS87] E. Olusegun George and Karan P. Singh. An approximation of F distribution by binomial probabilities. Statistics & Probability Letters, 5(3):169–173, April 1987. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0167715287900332.

Gupta:1982:TSD

[Gup82] Ashis Sen Gupta. Tests for simultaneously determining numbers of clusters and their shape with multivariate data. Statistics & Probability Letters, 1(1):46–50, July 1982. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016771528290013X.

Gupta:1983:UEF

[Gup83] R. P. Gupta. Unbiased estimation of finite population variance using auxiliary information. Statistics & Probability Letters, 1 (3):121–124, March 1983. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0167715283900585.

Ghahramani:1989:FMR

[GW89] Saeed Ghahramani and Ronald W. Wolff. Finiteness of moments of randomly stopped sums of i.i.d. random variables. Statistics & Probability Letters, 8(1):67–68, May 1989. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715289900850.

Gzyl:1987:CVV

[Gzy87] Henryk Gzyl. Characterization of vector valued, Gaussian, stationary, Markov processes. Statistics & Probability Letters, 6(1):17– 19, September 1987. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715287900538.

Hirano:1987:PED

[HA87] K. Hirano and S. Aki. Properties of the extended distributions of order κ. Statistics & Probability Letters, 6(2):67–69, November 1987. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715287900757.

Hart:1984:MDF

[Har84a] Jeffrey D. Hart. On the marginal distribution of a first order autoregressive process. Statistics & Probability Letters, 2(2):105– 109, March 1984. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715284900580.

Hart:1984:MRK

[Har84b] Jeffrey D. Hart. On the modal resolution of kernel density estimators. Statistics & Probability Letters, 2(6):363–369, December 1984. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715284900488.

Harris:1989:PLG

[Har89] Bernard Harris. Poisson limits for generalized random allocation problems. Statistics & Probability Letters, 8(2):123– 127, June 1989. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715289900047.

Hathaway:1986:AIE

[Hat86] Richard J. Hathaway. Another interpretation of the EM algorithm for mixture distributions. Statistics & Probability Letters, 4(2): 53–56, March 1986. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715286900167.

Huang:1989:NCG

[HC89] Wen-Jang Huang and Li-Sue Chen. Note on a characterization of gamma distributions. Statistics & Probability Letters, 8(5):485– 487, October 1989. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016771528990031X.

Hedayat:1987:SOM

[Hed87] A. Hedayat. On a statistical optimality of magic squares. Statistics & Probability Letters, 5(3):191–192, April 1987. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016771528790037X.

Herz:1988:SIO

[Her88] Carl Herz. Splitting intervals. Statistics & Probability Letters, 7 (1):3–7, July 1–87, 1988. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0167715288900788.

Hall:1988:LDE

[HH88] W. J. Hall and Wei-Min Huang. Large deviations and estimation in infinite-dimensional models. Statistics & Probability Letters, 6(6):433–439, May 1988. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0167715288901046.

Hallin:1983:NYW

[HI83] Marc Hallin and Jean-François Ingenbleek. Nonstationary Yule– Walker equations. Statistics & Probability Letters, 1(4):189– 195, June 1983. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715283900299.

Hahn:1988:UAN

[HK88] M. G. Hahn and J. Kuelbs. Universal asymptotic normality for conditionally trimmed sums. Statistics & Probability Letters, 7 (1):9–15, July 1–87, 1988. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.sciencedirect.

com/science/article/pii/016771528890079X.

Hall:1987:EIS

[HM87a] Peter Hall and J. S. Marron. Estimation of integrated squared density derivatives. Statistics & Probability Letters, 6(2):109–115, November 1987. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715287900836.

Huda:1987:MSO

[HM87b] S. Huda and Rahul Mukerjee. Minimax second-order designs for difference between estimated responses in extrapolation region. Statistics & Probability Letters, 6(1):43–45, September 1987. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715287900575.

Hall:1989:NAB

[HM89] Peter Hall and Michael A. Martin. A note on the accuracy of bootstrap percentile method confidence intervals for a quantile. Statistics & Probability Letters, 8(3):197–200, August 1989. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715289901211.

Hognas:1987:NPR

[Hög87] Göran Högnäs. A note on products of random matrices. Statistics & Probability Letters, 5(5):367–370, August 1987. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715287900125.

Holmquist:1987:ULR

[Hol87] Björn Holmquist. An unbiased likelihood ratio test for equality of the covariance matrices in several multivariate normal populations with partially known means. Statistics & Probability Letters, 5(2):99–103, March 1987. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0167715287900630.

Holst:1989:NBM

[Hol89] Lars Holst. A note on Banach’s match box problem. Statistics & Probability Letters, 8(5):441–443, October 1989. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715289900242.

Hoover:1989:BEO

[Hoo89] Donald R. Hoover. Bounds on expectations of order statistics for dependent samples. Statistics & Probability Letters, 8(3):261– 265, August 1989. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715289901314.

Hoppe:1985:IBB

[Hop85] Fred M. Hoppe. Iterating Bonferroni bounds. Statistics & Probability Letters, 3(3):121–125, June 1985. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715285900483.

Horvath:1984:SAC

[Hor84] Lajos Horváth. Strong approximation of certain stopped sums. Statistics & Probability Letters, 2(3):181–185, May 1984. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0167715284900117.

Horvath:1985:AAS

[Hor85] Lajos Horváth. Approximation for Abel sums of independent, identically distributed random variables. Statistics & Probability Letters, 3(4):221–225, July 1985. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0167715285900227.

Hougaard:1985:SAC

[Hou85] Philip Hougaard. Saddlepoint approximations for curved exponential families. Statistics & Probability Letters, 3(3):161– 166, June 1985. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715285900550.

Howard:1983:RSB

[How83] J. V. Howard. Random sequences of binary digits in which missing values can almost certainly be restored. Statistics & Probability Letters, 1(5):233–238, August 1983. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0167715283900378.

Hernandez:1987:THS

[HR87] Victor Hernández and Juan J. Romo. On the type hypothesis for the Strong Law of Large Numbers. Statistics & Probability Letters, 5(3):193–195, April 1987. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0167715287900381.

Husler:1989:CEN

[HR89a] J. Hüsler and H. Riedwyl. Comparison of the efficiency of nonparametric location tests based on grouped and individual data. Statistics & Probability Letters, 7(4):287–291, February 1989. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715289901077.

Husler:1989:MNR

[HR89b] Jürg Hüsler and Rolf-Dieter Reiss. Maxima of normal random vectors: Between independence and complete dependence. Statistics & Probability Letters, 7(4):283–286, February 1989. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715289901065.

Hettmansperger:1986:CIB

[HS86] Thomas P. Hettmansperger and Simon J. Sheather. Confidence intervals based on interpolated order statistics. Statistics & Probability Letters, 4(2):75–79, March 1986. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0167715286900210.

Hall:1989:LCV

[HS89] Peter Hall and William R. Schucany. A local cross-validation algorithm. Statistics & Probability Letters, 8(2):109–117, June 1989. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715289900023.

Hsing:1989:LMP

[Hsi89] Tailen Hsing. On a loss of memory property of the maximum. Statistics & Probability Letters, 8(5):493–496, October 1989. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715289900333.

Hung:1985:RET

[Hun85] Hsien-Ming Hung. Regression estimation with transformed auxiliary variates. Statistics & Probability Letters, 3(5):239–243, September 1985. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715285900380.

Hung:1986:PMF

[Hun86] Hsien-Ming Hung. On the prediction of means in a finite population. Statistics & Probability Letters, 4(3):113–118, April 1986. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715286900726.

Hung:1987:NEL

[Hun87] Hsien-Ming Hung. A note on extended least squares methods for cluster sampling. Statistics & Probability Letters, 5(5):343– 346, August 1987. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715287900083.

Husler:1985:GRP

[Hüs85] Jürg Hüsler. The growth rate of partial maxima on subsequences. Statistics & Probability Letters, 3(5):245–249, September 1985. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715285900392.

Huse:1988:APS

[Hus88] Vera R. Huse. Asymptotic properties for the sequential cusum procedure. Statistics & Probability Letters, 7(1):73–80, July 1– 87, 1988. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715288900910.

Hamedani:1985:PSR

[HW85] G. G. Hamedani and G. G. Walter. On the product of symmetric random variables. Statistics & Probability Letters, 3(5):251– 253, September 1985. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715285900409. See comments [HW86].

Hamedani:1986:PSR

[HW86] G. G. Hamedani and G. G. Walter. To: On the product of symmetric random variables. Statistics & Probability Letters, 4(1): 51, January 1986. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715286900404. See [HW85].

Hall:1988:MAD

[HW88] Peter Hall and Matthew P. Wand. On the minimization of absolute distance in kernel density estimation. Statistics & Probability Letters, 6(5):311–314, April 1988. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0167715288900053.

Iachan:1985:PSO

[Iac85] Ronaldo Iachan. Plane sampling. Statistics & Probability Letters, 3(3):151–159, June 1985. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0167715285900549.

Ihara:1987:SIS

[Iha87] Masamori Ihara. The structure of improper solutions in maximum likelihood factor analysis. Statistics & Probability Letters, 5(1):35–37, January 1987. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.sciencedirect.

com/science/article/pii/016771528790023X.

Kim:1987:UMEa

[iKB87a] Seong in Kim and D. S. Bai. On unbiased multinomial estimation. Statistics & Probability Letters, 5(1):29–34, January 1987. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715287900228.

Kim:1987:UMEb

[iKB87b] Seong in Kim and D. S. Bai. On unbiased multinomial estimation. Statistics & Probability Letters, 5(3):225–230, April 1987. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715287900459.

Ihara:1985:ECL

[IO85] Masamori Ihara and Masashi Okamoto. Experimental comparison of least-squares and maximum likelihood methods in factor analysis. Statistics & Probability Letters, 3(6):287–293, October 1985. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715285900574.

Ioannides:1987:NUC

[IR87] D. Ioannides and G. G. Roussas. Note on the uniform convergence of density estimates for mixing random variables. Statistics & Probability Letters, 5(4):279–285, June 1987. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715287901052.

Yoshihara:1987:CLT

[iYT87] Ken ichi Yoshihara and Hiroshi Takahata. The Central Limit Theorem for summability methods of some weakly dependent sequences. Statistics & Probability Letters, 5(2):143–147, March 1987. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715287900708.

Johnson:1989:NBA

[JC89] Wesley Johnson and Ronald Christensen. Nonparametric Bayesian analysis of the accelerated failure time model. Statistics & Probability Letters, 8(2):179–184, June 1989. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016771528990014X.

Jeyaratnam:1985:MVC

[Jey85] S. Jeyaratnam. Minimum volume confidence regions. Statistics & Probability Letters, 3(6):307–308, October 1985. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715285900616.

Johnson:1984:ALT

[JH84] Richard A. Johnson and James H. Haskell. An approximate lower tolerance bound for the three-parameter Weibull applied to lumber property characterization. Statistics & Probability Letters, 2(2): 67–76, March 1984. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715284900531.

Jammalamadaka:1984:LDN

[JJ84] S. Rao Jammalamadaka and Svante Janson. On the limiting distribution of the number of ‘near-matches’. Statistics & Probability Letters, 2(6):353–355, December 1984. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715284900464.

Jensen:1987:EPC

[JJ87] Søren Tolver Jensen and Søren Johansen. Estimation of proportional covariances. Statistics & Probability Letters, 6(2):83– 85, November 1987. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715287900782.

Jureckova:1988:MCT

[JKV88] J. Jurecková, W. C. M. Kallenberg, and N. Veraverbeke. Moderate and Cramér-type large deviation theorems for M-estimators. Statistics & Probability Letters, 6(3):191–199, February 1988. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0167715288901198.

Johnson:1982:P

[Joh82] Richard A. Johnson. Preface. Statistics & Probability Letters, 1 (1):1, July 1982. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715282900025.

John:1985:CBD

[Joh85] Peter W. M. John. Constructing balanced designs from triangular designs. Statistics & Probability Letters, 3(1):19–20, February 1985. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715285900057.

Johnson:1988:SPC

[Joh88] Dudley Paul Johnson. A short proof of the Central Limit Theorem for Markov chains. Statistics & Probability Letters, 6(4):225– 227, March 1988. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715288900648.

Jones:1987:IFM

[Jon87a] M. C. Jones. Inverse factorial moments. Statistics & Probability Letters, 6(1):37–42, September 1987. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715287900563.

Jones:1987:MRQ

[Jon87b] M. C. Jones. On moments of ratios of quadratic forms in normal variables. Statistics & Probability Letters, 6(2):129–136, November 1987. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715287900861.

Joshi:1987:NRG

[Jos87] V. M. Joshi. A note on resolution of Godambe’s paradox. Statistics & Probability Letters, 5(4):241, June 1987. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715287900976.

Jockel:1984:CVE

[JP84] Karl-Heinz Jöckel and Peter Pflaumer. Calculating the variance of errors in population forecasting. Statistics & Probability Letters, 2(4):211–213, August 1984. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0167715284900178.

Jennrich:1988:RDP

[JP88] Robert I. Jennrich and Sidney C. Port. Radial and directional parts of a random vector. Statistics & Probability Letters, 6(3): 155–158, February 1988. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0167715288901125.

Jensen:1986:NML

[JR86] Ernst Lykke Jensen and Holger Rootzén. A note on De Moivre’s limit theorems: Easy proofs. Statistics & Probability Letters, 4 (5):231–232, August 1986. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0167715286900945.

Kageyama:1985:SCS

[Kag85] Sanpei Kageyama. A structural classification of SR group divisible designs. Statistics & Probability Letters, 3(1):25–27, February 1985. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715285900070.

Kageyama:1988:EVB

[Kag88] Sanpei Kageyama. Existence of variance-balanced binary designs with fewer experimental units. Statistics & Probability Letters, 7 (1):27–28, July 1–87, 1988. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0167715288900831.

Kaminsky:1982:CPE

[Kam82] Kenneth S. Kaminsky. A characteristic property of the exponential distribution. Statistics & Probability Letters, 1(1):31– 32, July 1982. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715282900098.

Kano:1984:CAV

[Kan84] Yutaka Kano. Construction of additional variables conforming to a common factor model. Statistics & Probability Letters, 2(4):241– 244, August 1984. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715284900221.

Kao:1985:IDC

[Kao85] Chihwa Kao. Influence diagnostic for censored regression models. Statistics & Probability Letters, 3(6):337–342, October 1985. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715285900677.

Kappenman:1988:SMC

[Kap88] Russell F. Kappenman. A simple method for choosing between the lognormal and Weibull models. Statistics & Probability Letters, 7 (2):123–126, September 1988. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0167715288900363.

Katulska:1989:OBS

[Kat89] Krystyna Katulska. Optimum biased spring balance weighing designs. Statistics & Probability Letters, 8(3):267–271, August 1989. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715289901326.

Kochar:1985:ESS

[KD85] Subhash C. Kochar and Jayant V. Deshpande. On exponential scores statistics for testing against positive aging. Statistics & Probability Letters, 3(2):71–73, April 1985. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715285900276.

Keating:1984:NEP

[Kea84] Jerome P. Keating. A note on estimation of percentiles and reliability in the extreme-value distribution. Statistics & Probability Letters, 2(3):143–146, May 1984. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0167715284900051.

Kemp:1988:NSE

[Kem88] Adrienne W. Kemp. A note on Stirling’s expansion for factorial n. Statistics & Probability Letters, 7(2):139–143, September 1988. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715288900405.

Kettl:1989:NAA

[Ket89] Stena Kettl. A note on the ace algorithm. Statistics & Probability Letters, 8(1):27–28, May 1989. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0167715289900783.

Khalil:1986:CAD

[Kha86] Z. Khalil. On a class of asymptotic distributions in redundancy with renewals. Statistics & Probability Letters, 4(3):151– 155, April 1986. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715286900799.

Khuri:1987:ETN

[Khu87] A. I. Khuri. An exact test for the nesting effects variance component in an unbalanced random two-fold nested model. Statistics & Probability Letters, 5(4):305–311, June 1987. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715287901106.

Kindermann:1985:CFA

[KL85a] Ross P. Kindermann and Vincent N. LaRiccia. Closed form asymptotically efficient estimators based upon order statistics. Statistics & Probability Letters, 3(1):29–34, February 1985. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715285900082.

Koski:1985:ASI

[KL85b] Timo Koski and Wilfried Loges. Asymptotic statistical inference for a stochastic heat flow problem. Statistics & Probability Letters, 3(4):185–189, July 1985. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.sciencedirect.

com/science/article/pii/016771528590015X.

Klebaner:1986:RMG

[Kle86] Fima C. Klebaner. On the renewal measure for Gaussian sequences. Statistics & Probability Letters, 4(4):167–171, June 1986. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771528690060X.

Khatri:1987:CNW

[KM87a] C. G. Khatri and Rahul Mukerjee. Characterization of normality within the class of elliptical contoured distributions. Statistics & Probability Letters, 5(3):187–190, April 1987. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715287900368.

Kowalczyk:1987:SMN

[KM87b] Teresa Kowalczyk and Jan Mielniczuk. A screening method for a nonparametric model. Statistics & Probability Letters, 5(3):163– 167, April 1987. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715287900320.

Knight:1989:NAC

[Kni89] Keith Knight. A note on the asymptotic covariance matrix of the Yule–Walker estimator. Statistics & Probability Letters, 8(5):407– 410, October 1989. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715289900199.

Kocherlakota:1989:NBB

[Koc89] S. Kocherlakota. A note on the bivariate binomial distribution. Statistics & Probability Letters, 8(1):21–24, May 1989. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
016771528990076X.

Koenker:1984:NEL

[Koe84] Roger Koenker. A note on L-estimates for linear models. Statistics & Probability Letters, 2(6):323–325, December 1984. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0167715284900403.

Kokic:1987:RCS

[Kok87] P. N. Kokic. Rates of convergence in the Strong Law of Large Numbers for degenerate U-statistics. Statistics & Probability Letters, 5(5):371–374, August 1987. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0167715287900137.

Korwar:1985:JCD

[Kor85] R. M. Korwar. On the joint characterization of the distributions of the true income and the proportion of the reported income to the true income through a model of under-reported incomes. Statistics & Probability Letters, 3(4):201–203, July 1985. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715285900185.

Koronacki:1987:SAC

[Kor87a] Jacek Koronacki. A stochastic approximation counterpart of the feasible direction method. Statistics & Probability Letters, 5(6):415– 419, October 1987. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715287900927.

Korwar:1987:NEB

[Kor87b] Ramesh M. Korwar. Nonparametric estimation of a bivariate survivorship function with doubly censored data. Statistics & Probability Letters, 5(2):119–124, March 1987. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715287900666.

Korzeniowski:1989:DCE

[Kor89] Andrzej Korzeniowski. On diffusions that cannot escape from a convex set. Statistics & Probability Letters, 8(3):229–234, August 1989. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715289901272.

Koul:1985:MDE

[Kou85] Hira L. Koul. Minimum distance estimation in linear regression with unknown error distributions. Statistics & Probability Letters, 3(1):1–8, February 1985. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0167715285900021.

Kourouklis:1987:SCS

[Kou87] Stavros Kourouklis. On the strong consistency of a solution to the likelihood equation. Statistics & Probability Letters, 5(1):23– 27, January 1987. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715287900216.

Koziol:1985:NTS

[Koz85] James A. Koziol. A note on testing symmetry with estimated parameters. Statistics & Probability Letters, 3(4):227– 230, July 1985. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715285900239.

Kotz:1984:EEA

[KPW84] Samuel Kotz, W. L. Pearn, and Dean W. Wichern. Eigenvalueeigenvector analysis for a class of patterned correlation matrices with an application. Statistics & Probability Letters, 2(3):119– 125, May 1984. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715284900014.

Kremers:1988:EII

[Kre88] Walter Kremers. An extension and implications of the inspection paradox. Statistics & Probability Letters, 6(4):269– 273, March 1988. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715288900739.

Krzysko:1983:ADD

[Krz83] Miroslaw Krzyśko. Asymptotic distribution of the discriminant function. Statistics & Probability Letters, 1(5):243–250, August 1983. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715283900391.

Koul:1983:ESP

[KS83] H. L. Koul and V. Susarla. Estimators of scale parameters in linear regression. Statistics & Probability Letters, 1(5):273– 277, August 1983. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715283900445.

Koul:1985:KST

[KS85] H. L. Koul and P. K. Sen. On a Kolmogorov–Smirnov type aligned test in linear regression. Statistics & Probability Letters, 3(3): 111–115, June 1985. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016771528590046X.

Karatzas:1987:DBP

[KS87a] Ioannis Karatzas and Steven E. Shreve. A decomposition of the Brownian path. Statistics & Probability Letters, 5(2):87– 93, March 1987. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715287900617.

Kumar:1987:CFD

[KS87b] Srivastava Deo Kumar and M. Sreehari. Characterization of a family of discrete distributions via a Chernoff type inequality. Statistics & Probability Letters, 5(4):293–294, June 1987. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715287901076.

Kotz:1988:NCE

[KS88] S. Kotz and F. W. Steutel. Note on a characterization of exponential distributions. Statistics & Probability Letters, 6(3):201– 203, February 1988. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715288901204.

Kelly:1989:MPI

[KS89] Douglas Kelly and Gordon Simons. Markov processes with infinitely divisible limit distributions: some examples. Statistics & Probability Letters, 8(4):363–369, September 1989. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016771528990045X.

Kshirsagar:1989:EMO

[Ksh89] Anant M. Kshirsagar. Effect of a missing observation on Hotelling’s generalized T_{0}^{2}. Statistics & Probability Letters, 8(2):163–166, June 1989. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715289900102.

Kubokawa:1989:IEC

[Kub89] Tatsuya Kubokawa. Improved estimation of a covariance matrix under quadratic loss. Statistics & Probability Letters, 8(1): 69–71, May 1989. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715289900862.

Kuo:1986:NBE

[Kuo86] Lynn Kuo. A note on Bayes empirical Bayes estimation by means of Dirichlet processes. Statistics & Probability Letters, 4(3):145– 150, April 1986. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715286900787.

Kim:1985:BHD

[KV85] B. K. Kim and J. Van Ryzin. A bivariate histogram density estimator: Consistency and asymptotic normality. Statistics & Probability Letters, 3(3):167–173, June 1985. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715285900562.

Lariccia:1984:OWG

[Lar84] Vincent N. Lariccia. Optimal weights for general L^{2} distance estimators. Statistics & Probability Letters, 2(3):169–173, May 1984. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715284900099.

Linssen:1983:ERC

[LB83] H. N. Linssen and P. J. A. Banens. Estimation of the radius of a circle when the coordinates of a number of points on its circumference are observed: An example of bootstrapping. Statistics & Probability Letters, 1(6):307–311, October 1983. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715283900500.

Lusk:1982:UPA

[LC82] E. Lusk and B. Cassileth. The use of prediction analysis to develop groupings of survey questions. Statistics & Probability Letters, 1 (2):97–101, November 1982. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0167715282900232.

LeBreton:1988:NML

[Le 88] A. Le Breton. A note on maximum likelihood estimation for the complex-valued first-order autoregressive process. Statistics & Probability Letters, 7(2):171–173, September 1988. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715288900454.

Luxmann-Ellinghaus:1987:IMI

[LE87] U. Lüxmann-Ellinghaus. On the identifiability of mixtures of infinitely divisible power series distributions. Statistics & Probability Letters, 5(5):375–378, August 1987. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715287900149.

Lecoutre:1985:OCW

[Lec85] Jean-Pierre Lecoutre. The L_{2}-optimal cell width for the histogram. Statistics & Probability Letters, 3(6):303–306, October 1985. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715285900604.

Ledwina:1987:EIL

[Led87] Teresa Ledwina. An expansion of the index of large deviations for linear rank statistics. Statistics & Probability Letters, 5(4): 247–248, June 1987. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016771528790099X.

Lee:1988:SCP

[Lee88] Mei-Ling Ting Lee. Some cross-product difference statistics and a test for trends in ordered contingency tables. Statistics & Probability Letters, 7(1):41–46, July 1–87, 1988. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715288900867.

Lenk:1987:SRR

[Len87] Peter J. Lenk. Spectral representations for random densities. Statistics & Probability Letters, 5(4):295–298, June 1987. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0167715287901088.

Leskow:1989:NKS

[Leś89] Jacek Leśkow. A note of kernel smoothing of an estimator of a periodic function in the multiplicative intensity model. Statistics & Probability Letters, 7(5):395–400, April 1989. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016771528990093X.

Levin:1984:SSP

[Lev84] Bruce Levin. On a sequential selection procedure of Bechhofer, Kiefer, and Sobel. Statistics & Probability Letters, 2(2):91– 94, March 1984. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715284900567.

Lin:1989:VSM

[LH89] G. D. Lin and J. S. Huang. Variances of sample medians. Statistics & Probability Letters, 8(2):143–146, June 1989. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715289900072.

Li:1983:FME

[Li83] Tze Fen Li. A family of minimax estimators of a multivariate normal mean. Statistics & Probability Letters, 1(3):125– 128, March 1983. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715283900597.

Li:1984:FDM

[Li84] Tze Fen Li. A family of dominating minimax estimators of a multivariate normal mean. Statistics & Probability Letters, 2(4):215– 217, August 1984. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016771528490018X.

Li:1985:BEB

[Li85] Tze Fen Li. Bayes empirical Bayes estimation of a Poisson mean. Statistics & Probability Letters, 3(6):309–313, October 1985. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715285900628.

Li:1987:CTD

[Li87] Tze Fen Li. The cycle time distribution of exponential cyclic queues with multiple servers. Statistics & Probability Letters, 6 (2):77–81, November 1987. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0167715287900770.

Ling:1988:BDO

[Lin88] K. D. Ling. On binomial distributions of order k. Statistics & Probability Letters, 6(4):247–250, March 1988. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715288900697.

Ling:1989:NCN

[Lin89] K. D. Ling. A new class of negative binomial distributions of order k. Statistics & Probability Letters, 7(5):371–376, April 1989. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715289900540.

Liu:1986:ESF

[Liu86] Regina Y. Liu. Estimates for the score function under random censoring. Statistics & Probability Letters, 4(5):245–251, August 1986. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715286900970.

Lloyd:1987:OME

[Llo87] Christopher J. Lloyd. Optimal martingale estimating equations in a stochastic process. Statistics & Probability Letters, 5(6):381– 387, October 1987. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715287900873.

LeBreton:1987:SLL

[LM87] A. Le Breton and M. Musiela. A Strong Law of Large Numbers for vector Gaussian martingales and a statistical application in linear regression. Statistics & Probability Letters, 5(1):71– 73, January 1987. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715287900290.

Lio:1987:SCR

[LP87] Y. L. Lio and W. J. Padgett. Some convergence results for kerneltype quantile estimators under censoring. Statistics & Probability Letters, 5(1):5–14, January 1987. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715287900186.

Latour:1989:BSA

[LR89] Alain Latour and Roch Roy. On the behaviour of the sample autocovariances and autocorrelations of a seasonal Arima model. Statistics & Probability Letters, 8(4):339–345, September 1989. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715289900424.

Letac:1988:HDN

[LS88] Gérard Letac and V. Seshadri. Haight’s distributions as a natural exponential family. Statistics & Probability Letters, 6(3):165– 169, February 1988. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715288901150.

Luschgy:1989:CID

[Lus89] Harald Luschgy. Characterizations of infinite dimensional Gaussian shift experiments. Statistics & Probability Letters, 8(5):463– 468, October 1989. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016771528990028X.

Lusk:1982:NGS

[LW82] Eward J. Lusk and Haviland Wright. Non-Gaussian series and series with non-zero means: Practical implications for time series analysis. Statistics & Probability Letters, 1(1):2–6, July 1982. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715282900037.

Li:1989:ALA

[LY89] Ker-Chau Li and Donald Ylvisaker. Another look at adaptation on the average. Statistics & Probability Letters, 7(5):381– 383, April 1989. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715289900564.

Main:1988:ABR

[Mai88] P. Main. Asymptotic behaviour of reliability functions. Statistics & Probability Letters, 7(3):259–263, December 1988. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0167715288900600.

Marchetti:1983:AEE

[Mar83] Federico Marchetti. Asymptotic exponentiality of exit times. Statistics & Probability Letters, 1(4):167–170, June 1983. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
016771528390024X.

Maritz:1989:LEB

[Mar89] J. S. Maritz. Linear empirical Bayes estimation of quantiles. Statistics & Probability Letters, 8(1):59–65, May 1989. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715289900849.

Masry:1987:ASC

[Mas87] Elias Masry. Almost sure convergence of recursive density estimators for stationary mixing processes. Statistics & Probability Letters, 5(4):249–254, June 1987. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0167715287901003.

Matloff:1984:UCR

[Mat84] Norman S. Matloff. Use of covariates in randomized response settings. Statistics & Probability Letters, 2(1):31–34, January 1984. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715284900336.

Meijer:1987:EID

[MB87] J. W. Meijer and N. H. G. Baken. The exponential integral distribution. Statistics & Probability Letters, 5(3):209– 211, April 1987. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715287900411.

Meczarski:1985:SEM

[Mec85] Marek Meczarski. Sequential estimation of the minimum of a random variable. Statistics & Probability Letters, 3(2):63– 64, April 1985. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715285900252.

Meerschaert:1986:RVD

[Mee86] Mark M. Meerschaert. Regular variation and domains of attraction in R^{k}. Statistics & Probability Letters, 4(1):43–45, January 1986. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715286900386.

Muller-Funk:1989:HRC

[MFP89] Ulrich Müller-Funk and Friedrich Pukelsheim. How regular are conjugate exponential families? Statistics & Probability Letters, 7 (4):327–333, February 1989. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.sciencedirect.

com/science/article/pii/016771528990117X.

Mason:1985:SPC

[MG85] Robert L. Mason and Richard F. Gunst. Selecting principal components in regression. Statistics & Probability Letters, 3(6):299– 301, October 1985. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715285900598.

Mukerjee:1988:UOM

[MG88] Rahul Mukerjee and Sudhir Gupta. Universal optimality of main effect deletion designs. Statistics & Probability Letters, 7(2):89– 91, September 1988. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715288900314.

Mielniczuk:1987:ACB

[Mie87a] Jan Mielniczuk. Asymptotic confidence bands for densities based on nearest neighbor estimators under censoring. Statistics & Probability Letters, 5(2):125–128, March 1987. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715287900678.

Mielniczuk:1987:RCS

[Mie87b] Jan Mielniczuk. A remark concerning strong uniform consistency of the conditional Kaplan–Meier estimator. Statistics & Probability Letters, 5(5):333–337, August 1987. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016771528790006X.

Misiewicz:1989:PDN

[Mis89] Jolanta K. Misiewicz. Positive definite norm dependent functions on l^{∞}. Statistics & Probability Letters, 8(3):255–260, August 1989. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715289901302.

Mejza:1984:ISP

[MM84] Iwona Mejza and Stanislaw Mejza. Incomplete split plot designs. Statistics & Probability Letters, 2(6):327–332, December 1984. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715284900415.

McCune:1987:ICR

[MM87] E. D. McCune and S. K. McCune. On improving convergence rates for nonnegative kernel failure-rate function estimators. Statistics & Probability Letters, 6(2):71–76, November 1987. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715287900769.

Mack:1988:CTE

[MM88a] Y. P. Mack and Hans-Georg Müller. Convolution type estimators for nonparametric regression. Statistics & Probability Letters, 7(3):229– 239, December 1988. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715288900569. See erratum [Ano89d].

Mikulski:1988:ACR

[MM88b] Piotr W. Mikulski and Michael Monsour. On attainable Cramér– Rao type lower bounds for weighted loss functions. Statistics & Probability Letters, 7(1):1–2, July 1–87, 1988. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715288900776.

McCormick:1988:SCM

[MMR88] William P. McCormick, Ashim K. Mallik, and Jack H. Reeves. Strong consistency of the MLE for sequential design problems. Statistics & Probability Letters, 6(6):441–446, May 1988. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0167715288901058.

Marron:1988:CKD

[MN88] J. S. Marron and D. Nolan. Canonical kernels for density estimation. Statistics & Probability Letters, 7(3):195–199, December 1988. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715288900508.

Mori:1984:AIP

[MS84] Tamás F. Móri and Gábor J. Székely. Asymptotic independence of ‘pure head’ stopping times. Statistics & Probability Letters, 2(1): 5–8, January 1984. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715284900282.

Mazzuchi:1985:BAI

[MS85] Thomas A. Mazzuchi and Nozer D. Singpurwalla. A Bayesian approach to inference for monotone failure rates. Statistics & Probability Letters, 3(3):135–141, June 1985. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715285900513.

Mukherjea:1986:CED

[MS86] A. Mukherjea and D. Steele. Conditional expected durations of play given the ultimate outcome for a correlated random walk. Statistics & Probability Letters, 4(5):237–243, August 1986. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715286900969.

Maryak:1987:CIB

[MS87a] John L. Maryak and James C. Spall. Conditions for the insensitivity of the Bayesian posterior distribution to the choice of prior distribution. Statistics & Probability Letters, 5(6):401– 407, October 1987. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715287900903.

McKean:1987:CDL

[MS87b] Joseph W. McKean and Gerald L. Sievers. Coefficients of determination for least absolute deviation analysis. Statistics & Probability Letters, 5(1):49–54, January 1987. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715287900265.

Mukherjea:1987:OPC

[MS87c] Arunava Mukherjea and David A. Steele. Occupation probability of a correlated random walk and a correlated ruin problem. Statistics & Probability Letters, 5(2):105–111, March 1987. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715287900642.

Miceli:1988:AAM

[MS88] Robert Miceli and William E. Strawderman. “Almost arbitrary” minimax estimators of location for independent component variance mixtures of normal variates. Statistics & Probability Letters, 7 (2):131–133, September 1988. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0167715288900387.

Muller:1984:ODN

[Mül84] Hans-Georg Müller. Optimal designs for nonparametric kernel regression. Statistics & Probability Letters, 2(5):285–290, October 1984. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016771528490066X.

Mutafchiev:1988:LDN

[Mut88] Ljuben R. Mutafchiev. The limit distribution of the number of nodes in low strata of a random mapping. Statistics & Probability Letters, 7(3):247–251, December 1988. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0167715288900582.

Morrison:1987:CFS

[MW87] John M. Morrison and Gary L. Wise. Continuity of filtrations of sigma algebras. Statistics & Probability Letters, 6(1):55–60, September 1987. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715287900599.

Nayak:1988:NEN

[Nay88] Tapan K. Nayak. A note on estimating the number of errors in a system by recapture sampling. Statistics & Probability Letters, 7 (3):191–194, December 1988. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0167715288900491.

Nair:1989:CDM

[NH89] N. Unnikrishnan Nair and N. Hitha. Characterization of discrete models by distribution based on their partial sums. Statistics & Probability Letters, 8(4):335–337, September 1989. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0167715289900412.

Nickerson:1988:DPP

[Nic88] David M. Nickerson. Dominance of the positive-part version of the James–Stein estimator. Statistics & Probability Letters, 7(2): 97–103, September 1988. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0167715288900338.

Nagarsenker:1984:TET

[NN84] B. N. Nagarsenker and P. B. Nagarsenker. On a test of equality of two-parameter exponential distributions. Statistics & Probability Letters, 2(6):357–361, December 1984. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715284900476.

Nualart:1987:SRL

[Nua87] David Nualart. Some remarks on a linear stochastic differential equation. Statistics & Probability Letters, 5(3):231–234, April 1987. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715287900460.

Nelson:1988:SPK

[NY88] Paul I. Nelson and Shie-Shien Yang. Some properties of Kendall’s partial rank correlation coefficient. Statistics & Probability Letters, 6(3):147–150, February 1988. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0167715288901101.

OBrien:1982:DCI

[O’B82] G. L. O’Brien. Determining classes and independence. Statistics & Probability Letters, 1(2):57–59, November 1982. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715282900165.

Okamoto:1983:SMA

[OI83] Masashi Okamoto and Masamori Ihara. A simple Marquardt algorithm for the nonlinear least-squares problem. Statistics & Probability Letters, 1(6):301–305, October 1983. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715283900494.

Oja:1983:DSM

[Oja83] Hannu Oja. Descriptive statistics for multivariate distributions. Statistics & Probability Letters, 1(6):327–332, October 1983. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715283900548.

Oksanen:1987:NSU

[Oks87] E. H. Oksanen. A note on seemingly unrelated regression equations with residual vectors as explanatory variables. Statistics & Probability Letters, 6(2):103–105, November 1987. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715287900812.

ONeill:1985:TSI

[O’N85] Terence J. O’Neill. Testing for symmetry and independence in a bivariate exponential distribution. Statistics & Probability Letters, 3(5):269–274, September 1985. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0167715285900434.

Oneill:1986:IMP

[O’n86] Terence J. O’neill. Inconsistency of the misspecified proportional hazards model. Statistics & Probability Letters, 4(5):219– 222, August 1986. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016771528690091X.

OCinneide:1989:CME

[OR89] Colm A. O’Cinneide and Adrian E. Raftery. A continuous multivariate exponential distribution that is multivariate phase type. Statistics & Probability Letters, 7(4):323–325, February 1989. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715289901168.

Okazaki:1987:SKC

[OT87] Y. Okazaki and Y. Takahashi. Separating kernel of cylindrical measure. Statistics & Probability Letters, 5(6):397–399, October 1987. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715287900897.

Perez-Abreu:1988:ESS

[PA88] Victor Pérez-Abreu. The exponential space of an L^{2}-stochastic process with independent increments. Statistics & Probability Letters, 6(6):413–417, May 1988. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0167715288901010.

Papakonstantinou:1984:DLR

[Pap84] Vasilios Papakonstantinou. The distribution of length of N random unit vectors for a von Mises population. Statistics & Probability Letters, 2(2):111–115, March 1984. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0167715284900592.

Papathanasiou:1988:VBG

[Pap88] V. Papathanasiou. Variance bounds by a generalization of the Cauchy–Schwarz inequality. Statistics & Probability Letters, 7(1): 29–33, July 1–87, 1988. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0167715288900843.

Parzen:1982:MEI

[Par82] Emanuel Parzen. Maximum entropy interpretation of autoregressive spectral densities. Statistics & Probability Letters, 1(1): 7–11, July 1982. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715282900049.

Parr:1983:NHP

[Par83a] William C. Parr. A note on Hájek projections and the influence curve. Statistics & Probability Letters, 1(4):177–179, June 1983. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715283900263.

Parr:1983:IPD

[Par83b] William C. Parr. On the invariance principle for differentiable statistical functionals. Statistics & Probability Letters, 1(6):291– 293, October 1983. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715283900470.

Parr:1985:BSL

[Par85] William C. Parr. The bootstrap: Some large sample theory and connections with robustness. Statistics & Probability Letters, 3(2): 97–100, April 1985. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715285900331.

Park:1988:RSW

[Par88] Dong Ho Park. The reliability of k out of n systems when k components are equally reliable. Statistics & Probability Letters, 7 (3):175–178, December 1988. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0167715288900466.

Philippou:1988:MDO

[PAT88] Andreas N. Philippou, Demetris L. Antzoulakos, and Gregory A. Tripsiannis. Multivariate distributions of order κ. Statistics & Probability Letters, 7(3):207–216, December 1988. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715288900521.

Perruchet:1982:HCM

[Per82] Christophe Perruchet. Hierarchical classification of mathematical structures. Statistics & Probability Letters, 1(2):61–67, November 1982. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715282900177.

Pesotchinsky:1982:ECB

[Pes82] Leon Pesotchinsky. Expected cost bounds for the selection and ordering procedures based on binary-type questions. Statistics & Probability Letters, 1(1):41–45, July 1982. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715282900128.

Pesotchinsky:1984:MSO

[Pes84a] Leon Pesotchinsky. On a matching statistic. Statistics & Probability Letters, 2(5):305–310, October 1984. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715284900701.

Pesotchinsky:1984:ROP

[Pes84b] Leon Pesotchinsky. Robustness of the ordering procedures based on binary-type questions and on pairwise comparisons. Statistics & Probability Letters, 2(3):133–138, May 1984. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715284900038.

Peterson:1986:NSM

[Pet86] John J. Peterson. A note on some model selection criteria. Statistics & Probability Letters, 4(5):227–230, August 1986. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0167715286900933.

Peterson:1989:FSO

[Pet89] John J. Peterson. First and second order derivatives having applications to estimation of response surface optima. Statistics & Probability Letters, 8(1):29–34, May 1989. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715289900795.

Pflug:1983:KPC

[Pfl83] G. Ch. Pflug. On Kersting’s proof of the Central Limit Theorem. Statistics & Probability Letters, 1(6):323–326, October 1983. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715283900536.

Philippou:1983:GGD

[PGP83] Andreas N. Philippou, Costas Georghiou, and George N. Philippou. A generalized geometric distribution and some of its properties. Statistics & Probability Letters, 1(4):171–175, June 1983. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715283900251.

Phillips:1989:SMD

[Phi89] P. C. B. Phillips. Spherical matrix distributions and Cauchy quotients. Statistics & Probability Letters, 8(1):51–53, May 1989. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715289900825.

Piccioni:1984:PIL

[Pic84] Mauro Piccioni. A proof of the ineffectiveness of likelihood inference on singular measures. Statistics & Probability Letters, 2(2): 63–65, March 1984. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016771528490052X.

Piegorsch:1987:CBS

[Pie87] Walter W. Piegorsch. On confidence bands and set estimators for the simple linear model. Statistics & Probability Letters, 5 (6):409–413, October 1987. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0167715287900915.

Plachky:1988:CPC

[Pla88] D. Plachky. A continuity property of the convolution. Statistics & Probability Letters, 6(3):163–164, February 1988. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715288901149.

Philippou:1986:SRL

[PM86a] Andreas N. Philippou and Frosso S. Makri. Successes, runs and longest runs. Statistics & Probability Letters, 4(2):101– 105, March 1986. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715286900258.

Philippou:1986:SRLb

[PM86b] Andreas N. Philippou and Frosso S. Makri. Successes, runs and longest runs. Statistics & Probability Letters, 4(4):211– 215, June 1986. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715286900696.

Palachek:1983:CGR

[PS83] Albert D. Palachek and William R. Schucany. On the correlation of a group of rankings with an external ordering relative to the internal concordance. Statistics & Probability Letters, 1(5):259– 263, August 1983. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016771528390041X.

Pukelsheim:1986:IMW

[PT86] Friedrich Pukelsheim and D. Michael Titterington. Improving multiway block designs at the cost of nuisance parameters. Statistics & Probability Letters, 4(5):261–264, August 1986. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715286901008.

Padget:1988:SNQ

[PT88] W. J. Padget and L. A. Thombs. A smooth nonparametric quantile estimator from right-censored data. Statistics & Probability Letters, 7(2):113–121, September 1988. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0167715288900351.

Panaretos:1986:SGW

[PX86] John Panaretos and Evdokia Xekalaki. The stuttering generalized Waring distribution. Statistics & Probability Letters, 4(6):313– 318, October 1986. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715286900519.

Panaretos:1989:PDA

[PX89] John Panaretos and Evdokia Xekalaki. A probability distribution associated with events with multiple occurrences. Statistics & Probability Letters, 8(4):389–395, September 1989. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715289900497.

Rafajlowicz:1987:NOS

[Raf87] Ewaryst Rafajłowicz. Nonparametric orthogonal series estimators of regression: a class attaining the optimal convergence rate in L_{2}. Statistics & Probability Letters, 5(3):219–224, April 1987. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715287900447.

Rao:1984:ERC

[Rao84a] B. L. S. Prakasa Rao. On the exponential rate of convergence of the least squares estimator in the nonlinear regression model with Gaussian errors. Statistics & Probability Letters, 2(3):139– 142, May 1984. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016771528490004X.

Rao:1984:DEB

[Rao84b] P. V. Rao. Density estimation based on line transect samples. Statistics & Probability Letters, 2(1):51–57, January 1984. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0167715284900373.

Rao:1985:ATL

[Rao85] B. L. S. Prakasa Rao. Asymptotic theory of least squares estimator in a nonregular nonlinear regression model. Statistics & Probability Letters, 3(1):15–18, February 1985. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715285900045.

Rayner:1988:CUO

[Ray88] J. C. W. Rayner. Constructing a usable overlapping cells χ^{2} goodness of fit test. Statistics & Probability Letters, 6(4):257– 261, March 1988. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715288900715.

Rayner:1989:BIE

[Ray89] Robert K. Rayner. Bootstrap inversion of Edgeworth expansions for nonparametric confidence intervals. Statistics & Probability Letters, 8(3):201–206, August 1989. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0167715289901223.

Rodrigues:1987:KFM

[RB87] Josemar Rodrigues and Heleno Bolfarine. A Kalman filter model for single and two-stage repeated surveys. Statistics & Probability Letters, 5(4):299–303, June 1987. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.sciencedirect.

com/science/article/pii/016771528790109X.

Rigdon:1989:MSE

[RB89] Steven E. Rigdon and Asit P. Basu. Mean squared errors of estimators of the intensity function of a nonhomogeneous Poisson process. Statistics & Probability Letters, 8(5):445–449, October 1989. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715289900254.

Russo:1987:BCR

[RC87] Ralph P. Russo and Chern-Ching Chao. Boundary crossing random variables related to quantile convergence. Statistics & Probability Letters, 6(2):117–123, November 1987. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715287900848.

Rodrigues:1989:CFE

[RE89] Josemar Rodrigues and Silvia Nagib Elian. The coordinate-free estimation in finite population sampling. Statistics & Probability Letters, 7(4):293–295, February 1989. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715289901089.

Reiss:1986:NPA

[Rei86] Rolf-Dieter Reiss. A new proof of the approximate sufficiency of sparse order statistics. Statistics & Probability Letters, 4(5):233– 235, August 1986. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715286900957.

Rhee:1984:RCC

[Rhe84] WanSoo Rhee. On the rate of convergence in the Central Limit Theorem and the type of the Banach space. Statistics & Probability Letters, 2(2):59–62, March 1984. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0167715284900518.

Rhee:1986:CLT

[Rhe86] WanSoo T. Rhee. Central Limit Theorem and increment conditions. Statistics & Probability Letters, 4(4):191–195, June 1986. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715286900659.

Richards:1983:REE

[Ric83] Donald Richards. Representations for eigenfunctions of expected value operators in the Wishart distribution. Statistics & Probability Letters, 1(3):141–145, March 1983. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715283900627.

Richards:1985:PDS

[Ric85] Donald St. P. Richards. Positive definite symmetric functions on finite-dimensional spaces II. Statistics & Probability Letters, 3 (6):325–329, October 1985. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0167715285900653.

Rice:1986:CRP

[Ric86] John Rice. Convergence rates for partially splined models. Statistics & Probability Letters, 4(4):203–208, June 1986. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715286900672.

Rivest:1986:MKS

[Riv86] Louis-Paul Rivest. Modified Kent’s statistics for testing goodness of fit for the Fisher distribution in small concentrated samples. Statistics & Probability Letters, 4(1):1–4, January 1986. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715286900283.

Robbins:1983:NUP

[RL83] Herbert Robbins and Bruce Levin. A note on the ‘underadjustment phenomenon’. Statistics & Probability Letters, 1(3):137– 139, March 1983. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715283900615.

Rodrigues:1987:SSP

[Rod87] Josemar Rodrigues. Some shrunken predictors in finite populations with a multinormal superpopulation model. Statistics & Probability Letters, 5(5):347–351, August 1987. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715287900095.

Ronchetti:1985:RMS

[Ron85] Elvezio Ronchetti. Robust model selection in regression. Statistics & Probability Letters, 3(1):21–23, February 1985. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715285900069.

Rootzen:1984:ARC

[Roo84] Holger Rootzén. Attainable rates of convergence of maxima. Statistics & Probability Letters, 2(4):219–221, August 1984. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0167715284900191.

Rosalsky:1983:RFB

[Ros83] Andrew Rosalsky. A remark on the fluctuation behavior of I.I.D. Poisson random variables. Statistics & Probability Letters, 1(4): 181–182, June 1983. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715283900275. See acknowledgement

of priority [Ros84].

Rosalsky:1984:APR

[Ros84] Andrew Rosalsky. Acknowledgement of priority to “A remark on the fluctuation behavior of i.i.d. Poisson random variables”. Statistics & Probability Letters, 2(2):117, March 1984. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715284900609.

See [Ros83].

Rothe:1983:LBP

[Rot83] Günter Rothe. A lower bound for the Pitman efficiency of Friedman type tests. Statistics & Probability Letters, 1(5):239– 242, August 1983. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016771528390038X.

Rothe:1984:EPL

[Rot84] Günter Rothe. An efficiency property of linear nonparametric tests on trend. Statistics & Probability Letters, 2(4):197– 201, August 1984. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715284900142.

Roussas:1989:CRE

[Rou89a] George G. Roussas. Consistent regression estimation with fixed design points under dependence conditions. Statistics & Probability Letters, 8(1):41–50, May 1989. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0167715289900813.

Roussas:1989:SAP

[Rou89b] George G. Roussas. Some asymptotic properties of an estimate of the survival function under dependence conditions. Statistics & Probability Letters, 8(3):235–243, August 1989. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715289901284.

Rao:1985:MLT

[RP85] P. V. Rao and K. M. Portier. A model for line transect sampling clustered populations. Statistics & Probability Letters, 3(2): 89–93, April 1985. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715285900318.

Rubin:1983:CRL

[RR83] Herman Rubin and Andrew L. Rukhin. Convergence rates of large deviations probabilities for point estimators. Statistics & Probability Letters, 1(4):197–202, June 1983. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715283900305.

Rao:1986:ACM

[RS86a] B. L. S. Prakasa Rao and M. Sreehari. Another characterization of multivariate normal distribution. Statistics & Probability Letters, 4(4):209–210, June 1986. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0167715286900684.

Razanadrakoto:1986:DFM

[RS86b] Dieudonné Razanadrakoto and Norman C. Severo. The distribution of the ‘finitely many times’ in the Strong Law of Large Numbers. Statistics & Probability Letters, 4(4):179–182, June 1986. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715286900623.

Ramachandran:1988:PSR

[RS88] B. Ramachandran and V. Seshadri. On a property of strongly reproductive exponential families on R. Statistics & Probability Letters, 6(3):171–174, February 1988. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0167715288901162.

Rohatgi:1989:SIB

[RS89] Vijay K. Rohatgi and Gábor J. Székely. Sharp inequalities between skewness and kurtosis. Statistics & Probability Letters, 8 (4):297–299, September 1989. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0167715289900357.

Rao:1983:OTS

[RT83] J. S. Rao and Ram C. Tiwari. One- and two-sample match statistics. Statistics & Probability Letters, 1(3):129–135, March 1983. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715283900603.

Rhee:1986:MIJ

[RT86] WanSoo T. Rhee and Michel Talagrand. Martingale inequalities and the Jackknife estimate of variance. Statistics & Probability Letters, 4(1):5–6, January 1986. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0167715286900295.

Rudnicki:1984:LOR

[Rud84] Jerzy Rudnicki. Locally optimal rank tests for independence. Statistics & Probability Letters, 2(2):83–89, March 1984. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0167715284900555.

Rukhin:1983:CME

[Ruk83] Andrew L. Rukhin. A class of minimax estimators of a normal quantile. Statistics & Probability Letters, 1(5):217–221, August 1983. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715283900342.

Ruschendorf:1987:UEM

[Rüs87] Ludger Rüschendorf. Unbiased estimation of von Mises functionals. Statistics & Probability Letters, 5(4):287–292, June 1987. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715287901064.

Russo:1988:LET

[Rus88] Ralph P. Russo. On the last exit time and the number of exits of partial sums over a moving boundary. Statistics & Probability Letters, 6(6):371–377, May 1988. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0167715288900946.

Ruschendorf:1989:RTM

[Rüs89] Ludger Rüschendorf. On random translation models. Statistics & Probability Letters, 7(5):361–367, April 1989. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715289900527.

Rodriguez:1985:MEH

[RV85] Carlos C. Rodriguez and John Van Ryzin. Maximum entropy histograms. Statistics & Probability Letters, 3(3):117– 120, June 1985. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715285900471.

Robertson:1985:PIS

[RW85] James B. Robertson and James M. Womack. A pairwise independent stationary stochastic process. Statistics & Probability Letters, 3(4):195–199, July 1985. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0167715285900173.

Rychlik:1987:NDF

[Ryc87] Igor Rychlik. A note on Durbin’s formula for the first-passage density. Statistics & Probability Letters, 5(6):425–428, October 1987. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715287900940.

Samanta:1983:TIT

[Sam83] M. Samanta. On tests of independence in a trivariate exponential distribution. Statistics & Probability Letters, 1(6):279– 284, October 1983. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715283900457.

Sampson:1984:MCR

[Sam84] Allan R. Sampson. A multivariate correlation ratio. Statistics & Probability Letters, 2(2):77–81, March 1984. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715284900543.

Samanta:1989:NPE

[Sam89] M. Samanta. Non-parametric estimation of conditional quantiles. Statistics & Probability Letters, 7(5):407–412, April 1989. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0167715289900953.

Sandell:1988:GTP

[San88] Dennis Sandell. A game with three players. Statistics & Probability Letters, 7(1):61–63, July 1–87, 1988. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715288900892.

Sarma:1986:APM

[Sar86] Y. Ranakrishna Sarma. Asymptotic properties of maximum likelihood estimators from dependent observations. Statistics & Probability Letters, 4(6):309–311, October 1986. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715286900507.

Saleh:1984:TSU

[SAU84] A. K. MD. Ehsanes Saleh, M. Masoom Ali, and Dale Umbach. Tests of significance using selected sample quantiles. Statistics & Probability Letters, 2(5):295–297, October 1984. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715284900683.

Sawyer:1982:MDC

[Saw82] K. R. Sawyer. A multiple divergence criterion for testing between separate hypotheses. Statistics & Probability Letters, 1(1): 26–30, July 1982. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715282900086.

Shapiro:1988:ABE

[SB88] A. Shapiro and M. W. Browne. On the asymptotic bias of estimators under parameter drift. Statistics & Probability Letters, 7 (3):221–224, December 1988. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0167715288900545.

Scarsini:1985:LBD

[Sca85] Marco Scarsini. Lower bounds for the distribution function of a k-dimensional n-extensible exchangeable process. Statistics & Probability Letters, 3(2):57–62, April 1985. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715285900240.

Schott:1987:ICS

[Sch87] James R. Schott. An improved chi-squared test for a principal component. Statistics & Probability Letters, 5(5):361–365, August 1987. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715287900113.

Schott:1989:ATC

[Sch89a] James R. Schott. An adjustment for a test concerning a principal component subspace. Statistics & Probability Letters, 7(5): 425–430, April 1989. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715289900990.

Schucany:1989:LOW

[Sch89b] William R. Schucany. Locally optimal window widths for kernel density estimation with large samples. Statistics & Probability Letters, 7(5):401–405, April 1989. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0167715289900941.

Sen:1984:KST

[Sen84a] Pranab Kumar Sen. On a Kolmogorov–Smirnov type aligned test. Statistics & Probability Letters, 2(4):193–196, August 1984. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715284900130.

Seneta:1984:LSN

[Sen84b] E. Seneta. On the limiting set of nonnegative matrix products. Statistics & Probability Letters, 2(3):159–163, May 1984. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0167715284900075.

Sen:1987:WDA

[Sen87] Pranab Kumar Sen. What do the arithmetic, geometric and harmonic means tell us in length-biased sampling? Statistics & Probability Letters, 5(2):95–98, March 1987. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715287900629.

Serio:1984:TSS

[Ser84] Gabriella Serio. Two-stage stochastic model for carcinogenesis with time-dependent parameters. Statistics & Probability Letters, 2(2): 95–103, March 1984. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715284900579.

Seshadri:1988:SEI

[Ses88] V. Seshadri. A U-statistic and estimation for the inverse Gaussian distribution. Statistics & Probability Letters, 7(1):47–49, July 1–87, 1988. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715288900879.

Singh:1987:RRC

[SGS87] Gulab Singh, Sudhir Gupta, and Murari Singh. Robustness of row-column designs. Statistics & Probability Letters, 5(6):421– 424, October 1987. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715287900939.

Srivastava:1987:AMN

[SH87] M. S. Srivastava and T. K. Hui. On assessing multivariate normality based on Shapiro–Wilk W statistic. Statistics & Probability Letters, 5(1):15–18, January 1987. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0167715287900198.

Steutel:1989:HDB

[SH89] F. W. Steutel and B. G. Hansen. Haight’s distribution and busy periods. Statistics & Probability Letters, 7(4):301–302, February 1989. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715289901107.

Shao:1989:FCA

[Sha89a] Jun Shao. Functional calculus and asymptotic theory for statistical analysis. Statistics & Probability Letters, 8(5):397–405, October 1989. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715289900187.

Shao:1989:HSE

[Sha89b] Jun Shao. Half-sample estimation of sampling distributions. Statistics & Probability Letters, 8(2):147–155, June 1989. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0167715289900084.

Sheather:1986:FSE

[She86] Simon J. Sheather. A finite sample estimate of the variance of the sample median. Statistics & Probability Letters, 4(6):337– 342, October 1986. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715286900556.

Shi:1988:NDJ

[Shi88] Xiquan Shi. A note on the delete-d jackknife variance estimators. Statistics & Probability Letters, 6(5):341–347, April 1988. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715288900119.

Shih:1989:IIO

[Shi89] Weichung Joseph Shih. Influence of incomplete observations in multiple linear regression. Statistics & Probability Letters, 8(2): 171–174, June 1989. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715289900126.

Simons:1983:DAE

[Sim83] Gordon Simons. A discrete analogue and elementary derivation of ‘Lévy’s equivalence’ for Brownian motion. Statistics & Probability Letters, 1(4):203–206, June 1983. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715283900317.

Simons:1986:TVL

[Sim86] Gordon Simons. A trivariate version of ‘Lévy’s equivalence’. Statistics & Probability Letters, 4(1):7–8, January 1986. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0167715286900301.

Simonoff:1987:PES

[Sim87a] Jeffrey S. Simonoff. Probability estimation via smoothing in sparse contingency tables with ordered categories. Statistics & Probability Letters, 5(1):55–63, January 1987. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715287900277.

Simons:1987:EDW

[Sim87b] Gordon Simons. Easily determining which urns are ‘favorable’. Statistics & Probability Letters, 5(1):43–48, January 1987. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0167715287900253.

Sinha:1986:BER

[Sin86] S. K. Sinha. Bayesian estimation of the reliability function of the inverse Gaussian distribution. Statistics & Probability Letters, 4 (6):319–323, October 1986. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0167715286900520.

Singh:1987:MKE

[Sin87] Radhey S. Singh. Mise of kernel estimates of a density and its derivatives. Statistics & Probability Letters, 5(2):153–159, March 1987. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715287900721.

Sinha:1988:RHM

[Sin88] Kishore Sinha. Rectangular Hadamard matrices. Statistics & Probability Letters, 6(3):141–142, February 1988. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715288901083.

Sinha:1989:SNE

[Sin89] Kishore Sinha. Some new equireplicate balanced block designs. Statistics & Probability Letters, 8(1):89, May 1989. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0167715289900898.

Sinha:1988:FEB

[SJ88] K. Sinha and B. Jones. Further equireplicate balanced block designs with unequal block sizes. Statistics & Probability Letters, 6(4):229–230, March 1988. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.sciencedirect.

com/science/article/pii/016771528890065X.

Smythe:1986:UHC

[SKM86] Robert T. Smythe, Daniel Krewski, and Duncan Murdoch. The use of historical control information in modelling dose response relationships in carcinogenesis. Statistics & Probability Letters, 4(2): 87–93, March 1986. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715286900234.

Lau:1989:MI

[sL89] Tai shing Lau. Moment inequalities. Statistics & Probability Letters, 8(1):9–16, May 1989. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0167715289900746.

Szekely:1985:EPR

[SM85] G. J. Székely and T. F. Móri. An extremal property of rectangular distributions. Statistics & Probability Letters, 3(2):107– 109, April 1985. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715285900355.

Sheather:1987:CTC

[SM87] Simon J. Sheather and Joseph W. McKean. A comparison of testing and confidence interval methods for the median. Statistics & Probability Letters, 6(1):31–36, September 1987. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715287900551.

Smolenski:1985:AML

[Smo85] W. Smolenski. On the approximation of measurable linear functionals. Statistics & Probability Letters, 3(4):205–207, July 1985. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715285900197.

Seaman:1985:MVU

[SOY85] John W. Seaman, Jr., Patrick S. Odell, and Dean M. Young. Maximum variance unimodal distributions. Statistics & Probability Letters, 3(5):255–260, September 1985. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715285900410.

Smith:1984:LRT

[SR84] W. B. Smith and M. W. Riggs. Likelihood ratio testing on partial multinormal data. Statistics & Probability Letters, 2(6):337– 343, December 1984. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715284900439.

Srivastava:1984:MSK

[Sri84] M. S. Srivastava. A measure of skewness and kurtosis and a graphical method for assessing multivariate normality. Statistics & Probability Letters, 2(5):263–267, October 1984. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715284900622.

Sen:1983:SFM

[SS83] Pranab Kumar Sen and Ibrahim A. Salama. The Spearman footrule and a Markov chain property. Statistics & Probability Letters, 1 (6):285–289, October 1983. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0167715283900469.

Singh:1984:CEE

[SS84] R. Karan Singh and Gurdeep Singh. A class of estimators with estimated optimum values in sample surveys. Statistics & Probability Letters, 2(6):319–321, December 1984. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715284900397.

Sanz-Sole:1986:SRS

[SS86] Marta Sanz-Solé. Some remarks on stochastic differential equations in the plane with local Lipschitz coefficients. Statistics & Probability Letters, 4(6):343–348, October 1986. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715286900568.

Stadje:1986:EME

[Sta86] Wolfgang Stadje. An estimator for the mean of an exponential distribution with random right-censoring. Statistics & Probability Letters, 4(2):57–59, March 1986. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0167715286900179.

Stadje:1987:SME

[Sta87a] Wolfgan Stadje. On the SPRT for the mean of an exponential distribution. Statistics & Probability Letters, 5(6):389–395, October 1987. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715287900885.

Stadje:1987:NRP

[Sta87b] Wolfgang Stadje. Note on the renewal process for truncated exponential variables. Statistics & Probability Letters, 6(2):61–66, November 1987. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715287900745.

Stadje:1989:SDC

[Sta89] Wolfgang Stadje. Some distributions connected with sums of rectangular random variables. Statistics & Probability Letters, 8(2): 137–142, June 1989. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715289900060.

Stepniak:1986:AVC

[Ste86] Czeslaw Stepniak. On admissibility in various classes of quadratic estimators. Statistics & Probability Letters, 4(1):17–19, January 1986. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715286900325.

Steinebach:1988:OSA

[Ste88] Josef Steinebach. On the optimality of strong approximation rates for compound renewal processes. Statistics & Probability Letters, 6(4):263–267, March 1988. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0167715288900727.

Stigler:1982:PPD

[Sti82] Stephen M. Stigler. Poisson on the Poisson distribution. Statistics & Probability Letters, 1(1):33–35, July 1982. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715282900104.

Stout:1983:STU

[Sto83] William Stout. A statistical test of unidimensionality of a parameter underlying Bernoulli trails with applications. Statistics & Probability Letters, 1(4):183–188, June 1983. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715283900287.

Strittmatter:1989:FDP

[Str89] Wolfram Strittmatter. Finitely determined processes in metric spaces. Statistics & Probability Letters, 7(4):343–347, February 1989. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715289901193.

Sundberg:1985:WIR

[Sun85] Rolf Sundberg. When is the inverse regression estimator MSEsuperior to the standard regression estimator in multivariate controlled calibration situations? Statistics & Probability Letters, 3(2): 75–79, April 1985. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715285900288.

Svensson:1985:TGF

[Sve85] Åke Svensson. On χ^{2}-test of goodness-of-fit for a class of discrete multivariate models. Statistics & Probability Letters, 3(6): 331–336, October 1985. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715285900665.

Sweeting:1983:ISF

[Swe83] Trevor J. Sweeting. Independent scale-free spacings for the exponential and uniform distributions. Statistics & Probability Letters, 1(3):115–119, March 1983. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0167715283900573.

Seaman:1986:NIV

[SYM86] John W. Seaman, Jr., Dean M. Young, and Virgil R. Marco. A note on improved variance bounds for certain bounded unimodal distributions. Statistics & Probability Letters, 4(6):273– 274, October 1986. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715286900428.

Szekli:1988:NPS

[Sze88] R. Szekli. A note on preservation of self-decomposability under geometric compounding. Statistics & Probability Letters, 6(4):231– 236, March 1988. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715288900661.

Takahashi:1987:NCD

[Tak87] Rinya Takahashi. Normalizing constants of a distribution which belongs to the domain of attraction of the Gumbel distribution. Statistics & Probability Letters, 5(3):197–200, April 1987. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715287900393.

Tallis:1986:OBS

[Tal86] G. M. Tallis. On the optimality of balanced sampling. Statistics & Probability Letters, 4(3):141–144, April 1986. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715286900775.

Tan:1986:SGB

[Tan86] W. Y. Tan. A stochastic Gompertz birth-death process. Statistics & Probability Letters, 4(1):25–28, January 1986. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715286900349.

Taylor:1982:BFR

[Tay82] James R. Taylor. Backward and forward recurrence times, and their interrelationships in Bellman–Harris branching processes. Statistics & Probability Letters, 1(1):12–16, July 1982. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715282900050.

Ting:1989:OSC

[TBGG89] Naitee Ting, Richard K. Burdick, Franklin A. Graybill, and Rongde Gui. One-sided confidence intervals on nonnegative sums of variance components. Statistics & Probability Letters, 8(2):129– 135, June 1989. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715289900059.

Theil:1988:RBT

[TC88] Henri Theil and Ching-Fan Chung. Relations between two sets of variates: the bits of information provided by each variate in each set. Statistics & Probability Letters, 6(3):137–139, February 1988. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715288901071.

Tang:1984:DPI

[TG84] Jen Tang and A. K. Gupta. On the distribution of the product of independent beta random variables. Statistics & Probability Letters, 2(3):165–168, May 1984. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0167715284900087.

Tableman:1987:TSI

[TH87] Mara Tableman and Thomas P. Hettmansperger. Two-sample inference based on one-sample Wilcoxon signed rank statistics. Statistics & Probability Letters, 6(2):97–102, November 1987. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715287900800.

Theil:1982:SRN

[The82] Henri Theil. Some recent and new results on the maximum entropy distribution. Statistics & Probability Letters, 1(1):17– 22, July 1982. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715282900062.

Theil:1987:HMB

[The87] Henri Theil. How many bits of information does an independent variable yield in a multiple regression? Statistics & Probability Letters, 6(2):107–108, November 1987. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0167715287900824.

Hung:1984:NOG

[tHK84] Yung tai Hung and Anant M. Kshirsagar. A note on optimum grouping and the relative discriminating power of qualitative to continuous normal variates. Statistics & Probability Letters, 2(1): 19–21, January 1984. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715284900312.

Thompson:1989:TUB

[Tho89] Patrick A. Thompson. A transformation useful for bounding a forecast. Statistics & Probability Letters, 8(5):469–475, October 1989. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715289900291.

Too:1989:CUE

[TL89] Yeu-Hua Too and Gwo Dong Lin. Characterizations of uniform and exponential distributions. Statistics & Probability Letters, 7 (5):357–359, April 1989. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0167715289900515.

Theil:1984:CND

[TRF84] Henri Theil, Mercedes C. Rosalsky, and Renate Finke. A comparison of normal and discrete bootstraps for standard errors in equation systems. Statistics & Probability Letters, 2(3):175– 180, May 1984. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715284900105.

Titterington:1989:IMV

[TS89] D. M. Titterington and J. Sedransk. Imputation of missing values using density estimation. Statistics & Probability Letters, 8 (5):411–418, October 1989. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0167715289900205.

Tsai:1985:RTC

[Tsa85] Wei-Yann Tsai. Rank tests for a class of semi-Markov models with censored matched pairs. Statistics & Probability Letters, 3 (5):281–286, September 1985. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0167715285900458.

Tsai:1989:BNR

[Tsa89] Chih-Ling Tsai. Bias in nonlinear regression model with heteroscedastic or Ar(1) error structure. Statistics & Probability Letters, 8(2):167–170, June 1989. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0167715289900114.

Tsujitani:1988:GAR

[Tsu88a] Masaaki Tsujitani. Graphical analysis of residuals in stratified four-fold tables. Statistics & Probability Letters, 6(5):335– 339, April 1988. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715288900107.

Tsujitani:1988:OSA

[Tsu88b] Masaaki Tsujitani. Optimal scaling for association models when category scores have a natural ordering. Statistics & Probability Letters, 6(3):175–180, February 1988. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0167715288901174.

Tan:1989:SIB

[TX89] Xiaodong Tan and Yuan Xu. Some inequalities of Bonferroni– Galambos type. Statistics & Probability Letters, 8(1):17–20, May 1989. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715289900758.

Veraverbeke:1987:KTE

[Ver87] Noël Veraverbeke. A kernel-type estimator for generalized quantiles. Statistics & Probability Letters, 5(3):175–180, April 1987. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715287900344.

Vervaat:1988:NVC

[Ver88] Wim Vervaat. Narrow and vague convergence of set functions. Statistics & Probability Letters, 6(5):295–298, April 1988. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0167715288900028.

Vitale:1988:DVE

[Vit88] Richard A. Vitale. A differential version of the Efron–Stein inequality: bounding the variance of a function of an infinitely divisible variable. Statistics & Probability Letters, 7(2):105–112, September 1988. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016771528890034X.

Vardeman:1983:AEP

[VM83] Stephen Vardeman and Glen Meeden. Admissible estimators of the population total using trimming and Winsorization. Statistics & Probability Letters, 1(6):317–321, October 1983. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715283900524.

Wallenstein:1985:RMC

[Wal85] Sylvan Wallenstein. A regression model to combine results from 2 × 2 tables. Statistics & Probability Letters, 3(3):147– 150, June 1985. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715285900537.

Wansbeek:1985:EEA

[Wan85] Tom Wansbeek. Eigenvalue–eigenvector analysis for a class of patterned correlation matrices with an application: a comment. Statistics & Probability Letters, 3(2):95–96, April 1985. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
016771528590032X.

Williamson:1988:IDU

[WD88] R. C. Williamson and T. Downs. The inverse and determinant of a 2 × 2 uniformly distributed random matrix. Statistics & Probability Letters, 7(2):167–170, September 1988. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715288900442.

Weber:1984:RTR

[Web84] N. C. Weber. On resampling techniques for regression models. Statistics & Probability Letters, 2(5):275–278, October 1984. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715284900646.

Wefelmeyer:1985:CCM

[Wef85] Wolfgang Wefelmeyer. A counterexample concerning monotone unimodality. Statistics & Probability Letters, 3(2):87–88, April 1985. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715285900306.

Weiss:1986:NSS

[Wei86] Lionel Weiss. A note on small-sample maximum probability estimation. Statistics & Probability Letters, 4(3):109–111, April 1986. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715286900714.

Weiss:1987:ENM

[Wei87] Lionel Weiss. Estimating normal means with symmetric gain functions. Statistics & Probability Letters, 6(1):7–9, September 1987. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715287900502.

Welsh:1984:NSE

[Wel84] A. H. Welsh. A note on scale estimates based on the empirical characteristic function and their application to test for normality. Statistics & Probability Letters, 2(6):345–348, December 1984. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715284900440.

Welsh:1986:IEC

[Wel86] A. H. Welsh. Implementing empirical characteristic function procedures. Statistics & Probability Letters, 4(2):65–67, March 1986. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715286900192.

Welsh:1988:AEE

[Wel88] A. H. Welsh. Asymptotically efficient estimation of the sparsity function at a point. Statistics & Probability Letters, 6(6): 427–432, May 1988. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715288901034.

Wesolowski:1987:RCN

[Wes87] Jacek Wesołowski. A regressional characterization of the normal law. Statistics & Probability Letters, 6(1):11–12, September 1987. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715287900514.

Whitney:1987:NEI

[Whi87] Paul Whitney. A note on estimation and information topologies. Statistics & Probability Letters, 5(2):149–151, March 1987. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771528790071X.

Wiggins:1985:UDD

[Wig85] Alvin D. Wiggins. On the use of the directional derivative in obtaining multivariate extreme values. Statistics & Probability Letters, 3(1):39–44, February 1985. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0167715285900100.

Willmot:1988:RPP

[Wil88] Gordon E. Willmot. A remark on the Poisson–Pascal and some other contagious distributions. Statistics & Probability Letters, 7 (3):217–220, December 1988. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0167715288900533.

Wittmann:1985:ARM

[Wit85] Rainer Wittmann. An application of Rosenthal’s moment inequality to the Strong Law of Large Numbers. Statistics & Probability Letters, 3(3):131–133, June 1985. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0167715285900501.

Woodroofe:1988:STL

[WJ88] Michael Woodroofe and Myoungshic Jhun. Singh’s theorem in the lattice case. Statistics & Probability Letters, 7(3):201–205, December 1988. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016771528890051X.

Wansbeek:1983:NSD

[WK83] Tom Wansbeek and Arie Kapteyn. A note on spectral decomposition and maximum likelihood estimation in ANOVA models with balanced data. Statistics & Probability Letters, 1(4):213– 215, June 1983. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715283900330.

Wu:1989:BTP

[Wu89] Zizhi Wu. On Bayes’ test for the population mean of bounded observations. Statistics & Probability Letters, 7(5):385–389, April 1989. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715289900916.

Yandell:1988:BDS

[Yan88] Brian S. Yandell. Block diagonal smoothing splines. Statistics & Probability Letters, 6(5):331–334, April 1988. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715288900090.

Yannaros:1989:CRP

[Yan89] Nikos Yannaros. On Cox and renewal processes. Statistics & Probability Letters, 7(5):431–433, April 1989. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715289901004.

Yao:1988:ENC

[Yao88] Yi-Ching Yao. Estimating the number of change-points via Schwarz’ criterion. Statistics & Probability Letters, 6(3):181–189, February 1988. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715288901186.

Yao:1989:SCN

[Yao89] Y. C. Yao. A simple characterization of non-randomized admissible procedures in group testing. Statistics & Probability Letters, 8(2):119–122, June 1989. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0167715289900035.

Ycart:1989:MPE

[Yca89] Bernard Ycart. Markov processes and exponential families on a finite set. Statistics & Probability Letters, 8(4):371–376, September 1989. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715289900461.

Ying:1989:NAP

[Yin89] Zhiliang Ying. A note on the asymptotic properties of the productlimit estimator on the whole line. Statistics & Probability Letters, 7(4):311–314, February 1989. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0167715289901132.

Yang:1984:SIB

[YM84] Mark C. K. Yang and Tse Chin Mo. Some improvements on the Birnbaum–McCarty bound for P (Y < X). Statistics & Probability Letters, 2(3):127–132, May 1984. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715284900026.

Yeo:1989:CED

[YM89] G. F. Yeo and R. K. Milne. On characterizations of exponential distributions. Statistics & Probability Letters, 7(4):303–305, February 1989. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715289901119.

Yakovlev:1988:BCS

[YR88] A. Yu Yakovlev and S. T. Rachev. Bounds for crude survival probabilities within competing risks framework and their statistical application. Statistics & Probability Letters, 6(6):389– 394, May 1988. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715288900971.

Young:1987:NKI

[YSM87] Dean M. Young, John W. Seaman, Jr., and Virgil R. Marco. A note on a Kolmogorov inequality. Statistics & Probability Letters, 5(3):217–218, April 1987. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0167715287900435.

Zelterman:1986:LLR

[Zel86] D. Zelterman. The log-likelihood ratio for sparse multinomial mixtures. Statistics & Probability Letters, 4(2):95–99, March 1986. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715286900246.

Zelterman:1988:LRT

[Zel88] D. Zelterman. Likelihood ratio tests for central mixtures. Statistics & Probability Letters, 6(4):275–279, March 1988. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715288900740.

Zielinski:1986:QRN

[Zie86] Ryszard Zieliński. A quasi-random number generator with infinite period. Statistics & Probability Letters, 4(5):259, August 1986. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715286900994.

Zacks:1986:NMV

[ZR86] S. Zacks and Josemar Rodriguez. A note on the missing value principle and the EM-Algorithm for estimation and prediction in sampling from finite populations with a multinormal superpopulation model. Statistics & Probability Letters, 4(1):35–37, January 1986. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715286900362.